### v1

# Kilohertz Gravitational Waves from Binary Neutron Star Mergers: Full Spectrum Analyses and High-density Constraints on Neutron Star Matter

Abstract
We demonstrate Bayesian analyses of the complete gravitational-wave spectrum of binary neutron star mergers events with the next-generation detector Einstein Telescope. Our mock analyses are performed for 20 different signals using the TEOBResumSPA_NRPMw waveform that models gravitational-waves from the inspiral to the postmerger phase. They are employed to validate a pipeline for neutron star’s extreme matter constraints with prospective detections and under minimal hypotheses on the equation of state. The proposed analysis stack delivers inferences for the mass-radius curve, the mass dependence of the quadrupolar tidal polarizability parameter, the neutron star’s maximum density, the maximum mass and the relative radius, and the pressure-density relation itself. We show that a single event at a signal-to-noise ratio close to the minimum threshold for postmerger detection is sufficient to tightly constrain all the above relations as well as quantities like the maximum mass (maximum density) to precision of ${\sim}6$% (${\sim}10$%) at 90% credibility level. We also revisit inferences of prompt black hole formation with full spectrum signals and find that the latter can be robustly identified, even when the postmerger is not detectable due to a low signal-to-noise ratio. New results on the impact of the initial signal frequency and of the detector configuration (triangular vs. two-L) on the source’s parameters estimation are also reported. An improvement of approximately one order of magnitude in the precision of the chirp mass and mass ratio can be achieved by lowering the initial frequency from 20 Hz to 2 Hz. The two-L configuration shows instead significant improvements on the inference of the source declination, due to geographical separation of the two detectors.
pacs:
04.30.Db, 95.30.Sf, 97.60.Jd
I Introduction
Binary neutron stars are the main sources for next-generation (XG) ground-based gravitational wave (GW) detectors, such as Einstein Telescope (ET) Hild et al. (2011); Hild (2012); Abac et al. (2025); Maggiore et al. (2020); Punturo et al. (2010a) and Cosmic Explorer (CEReitze et al. (2019); Evans et al. (2021). They are expected to be observed with a rate of $7\times 10^{4}$ mergers per year Belgacem et al. (2019); Kalogera et al. (2021); Iacovelli et al. (2022). ET is designed with a xylophone configuration, combining a cryogenic interferometer for the low-frequency band with a high-power interferometer targeting frequencies above ${\sim}100~{}{\rm Hz}$. This design will not only significantly enhance current GW detector sensitivity, but also enable the observations of postmerger (PM) signals at kiloHertz frequencies. GWs observations of the full spectrum (inspiral-merger-postmerger) of a BNS merger can convey unique information on the nuclear matter that constitutes this compact object e.g. (Abbott et al., 2018a; Bauswein et al., 2017; Breschi et al., 2019; Radice et al., 2020; Breschi et al., 2022a; Wijngaarden et al., 2022), thus improving current GW constraints after GW170817 Abbott et al. (2017a, b, 2018b). While the inspiral-merger (IM) carries the footprint of the matter properties at the progenitor densities, i.e. $\rho\lesssim{2}\,\rho_{\rm sat}$ (where $\rho_{\rm sat}\simeq 2.7{\times}10^{14}~{}{\rm g}~{}{\rm cm}^{-3}$ is the nuclear saturation density), the kiloHertz PM transient encodes information about the high-density (${\sim}3-6\rho_{\rm sat}$) regime of the equation of state (EoSBreschi et al. (2022a), enabling the investigation of unrevealed properties of the nuclear matter at high-densities, unreachable with modern experiments Radice et al. (2017); Bauswein et al. (2019); Breschi et al. (2022a); Most and Raithel (2021); Prakash et al. (2021).
Inference of source properties and constraints on the neutron star (NS) EoS requires the availability of complete waveform templates for matched-filtering analyses. These templates are crucially designed using empirical relations computed from numerical relativity (NR) simulations and connecting morphological features of the PM signal to NS properties, e.g. Shibata and Uryu (2002); Stergioulas et al. (2011); Bauswein and Janka (2012); Bauswein et al. (2012); Hotokezaka et al. (2013); Takami et al. (2014); Bauswein and Stergioulas (2015); Bernuzzi et al. (2015); Breschi et al. (2022a, 2024a). In particular, Breschi et al. (2019) proposed the first time-domain full spectrum model that was constructed by combining an effective-one-body (EOB) IM waveform with a NR-informed analytical PM model (see also Breschi et al. (2024a, 2022b)). Similar approaches have been used in Wijngaarden et al. (2022); Puecher et al. (2023a), where phenomenological frequency-domain IM waveforms are stitched to PM models based on wavelets or Lorenzian templates. In all cases the phenomenological PM parameters are connected to the source parameter using NR empirical relations; see also (Bauswein et al., 2016; Bose et al., 2018; Soultanis et al., 2022; Vretinaris et al., 2025) for other examples of analytical PM models and (Easter et al., 2020; Whittaker et al., 2022) for data-driven PM models, which have not been incorporated in full spectrum models yet.
These modeling efforts, combined with the exceptional sensitivity of XG detectors, provide a crucial foundation for accurate EoS inference. Notably, even a single detection of BNS with ET may suffice to place stringent constraints on the NS mass-radius relation Breschi et al. (2022a). That approach enabled tight constraints of the mass-radius diagram and demonstrated that the maximum NS mass can be determined with an accuracy better than 12%. Additional works Breschi et al. (2023); Rezzolla and Takami (2016); Chatziioannou et al. (2017); Lehoucq et al. (2025); Vretinaris et al. (2025) have similarly combined information from the IM section to PM features, highlighting the complementarity of these two regimes in probing supranuclear matter. Numerous proof of principle studies focus exclusively on the PM GW emission from BNSs to investigate the high-density nuclear matter through NR simulations. These works exploit the connection between peak frequencies and EoS-dependent properties, such as total mass and compactness Bauswein et al. (2014); Most and Raithel (2021); Easter et al. (2019); Takami et al. (2015); Bauswein et al. (2016); Ecker et al. (2025); Bamber et al. (2025); Mitra et al. (2025). Furthermore, recent works explored the potential for EoS inference using advanced and XG detectors Walker et al. (2024); Gupta et al. (2022); Iacovelli et al. (2023); Bandopadhyay et al. (2024); Kiendrebeogo et al. (2023) and assess the impact of different waveform models on nuclear EoS inference Yelikar et al. (2024).
In the cases where the presence of the PM is not detectable, a major issue is the inference of prompt (rapid) black hole (BH) formation after merger. While electromagnetic counterparts can provide valuable information in this regard Abbott et al. (2017b); Coulter et al. (2017); Chornock et al. (2017); Nicholl et al. (2017); Cowperthwaite et al. (2017); Tanvir et al. (2017); Tanaka et al. (2017), GWs remain the key messenger to identify unambiguously the presence of a remnant BH. Many authors have explored the conditions under which a prompt collapse (PC) occurs instead of the formation of a NS remnant. This is primarily connected to the total mass of the BNS and the EoS of dense nuclear matter Shibata et al. (2005); Hotokezaka et al. (2011); Bauswein et al. (2013); Kashyap et al. (2022); Perego et al. (2022). In addition, PC analysis can be employed to place constraints on the EoS Bauswein et al. (2017, 2020); Tootle et al. (2021); Kashyap et al. (2022). Agathos et al. (2020) presented the first comprehensive Bayesian analysis to assess the probability of PC directly from the inspiral GW signal. This methodology was subsequently applied in the analysis of GW190425 Abbott et al. (2020a).
In this paper, we perform an extensive investigation of the capabilities of ET to infer BNS properties. We firstly focus on the low-frequency regime, assessing the impact on source parameter estimation of the initial frequency and of detector configurations (two-L and triangular). We then concentrate in the detections of the complete GW spectrum and study the prospective accuracy of EoS inference. We develop and employ a full spectrum model which combines the fast frequency-domain EOB TEOBResumSPA Akcay et al. (2019); Gamba et al. (2021a, 2023) with the frequency-domain wavelet-based NRPMw model Breschi et al. (2024a, 2022b) (Paper I and Paper II hereafter). We then perform Bayesian mock analyses of 20 fiducial BNSs, covering the frequency range from $f_{0}=20$ Hz up to the kiloHertz regime. These analyses serve not only to demonstrate the feasibility of combining information across the full GW spectrum, from inspiral to merger and postmerger, but also to quantify the achievable accuracy in the inference of source parameters. Beyond parameter estimation, our primary objective is to assess the improvements in constraints on the supranuclear properties governing NS matter. By incorporating both low- and high-frequency components of the signal, we explore how analyses of ET signals enhance our ability to probe the dense-matter regime, including tighter bounds on the mass-radius relation, the maximum mass supported by a stable non-rotating NS and the PC inference.
The paper is structured as follows. In Sec. II we introduce the full spectrum waveform model and the framework used for the analyses. We present the results in Sec. III, covering both inspiral only and full spectrum analyses, as well as a discussion about the EoS constraints and the probability of prompt BH formation. Finally Sec. IV provides a summary and our main conclusion. Moreover, we include Appendices about the waveform model and Bayesian analysis.
Conventions –
All quantities are expressed in SI units, with masses in Solar masses ${\rm M_{\odot}}$ and distances in Mpc. The symbol ‘$\log$’ denotes the natural logarithm. The waveform strain is decomposed in multipoles $(\ell,m)$ as
$h_{+}-{\rm i}h_{\times}=D_{L}^{-1}\sum_{\ell=2}^{\infty}\sum_{m=-\ell}^{\ell}h_{\ell m}(t)\,{}_{-2}Y_{\ell m}(\iota,\varphi),$ (1)
where $D_{L}$ is the luminosity distance, ${}_{-2}Y_{\ell m}$ are the $s=-2$ spin-weighted spherical harmonics and $\iota,\varphi$ are respectively the polar and azimuthal angles that define the orientation of the binary with respect to the observer. Each waveform mode $h_{\ell m}(t)$ is decomposed in amplitude $A_{\ell m}(t)$ and phase $\phi_{\ell m}(t)$, as
$h_{\ell m}(t)=A_{\ell m}(t)e^{-{\rm i}\phi_{\ell m}(t)}\,,$ (2)
with a related GW frequency,
$\omega_{\ell m}(t)=2\pi f_{\ell m}(t)=\frac{{\rm d}}{{\rm d}t}{\phi_{\ell m}}(t)\,.$ (3)
The moment of merger is defined as the time of the peak of $A_{22}(t)$. When the $(\ell,m)$ indices are omitted, we consider only the dominant $(2,2)$ mode.
The waveform depends on intrinsic and extrinsic source parameters. The intrinsic parameters of a BNS system are masses, spins and tidal polarizability parameters, e.g. $\boldsymbol{\theta}_{\rm bin}=\{M,q,\chi_{1},\chi_{2},\Lambda_{1},\Lambda_{2}\}$. The total binary mass is indicated with $M=m_{1}+m_{2}$, the mass ratio is $q=m_{1}/m_{2}\geq 1$, and the symmetric mass ratio is $\nu=m_{1}m_{2}/M^{2}$. The dimensionless spin vectors are denoted with ${\boldsymbol{\chi}}_{i}$ for $i=1,2$ and the spin component aligned with the orbital angular momentum L are labeled as $\chi_{i}={\boldsymbol{\chi}}_{i}\cdot\textbf{L}/|\textbf{L}|$. The effective spin parameter ${\chi_{\rm eff}}$ is a mass-weighted, aligned spin combination, i.e.
${\chi_{\rm eff}}=\frac{m_{1}\chi_{1}+m_{2}\chi_{2}}{M}\,.$ (4)
The quadrupolar tidal polarizability parameters are defined as $\Lambda_{i}=({2}/{3})\,k_{2,i}\,C_{i}^{-5}$ for $i=1,2$, where $k_{2,i}$ and $C_{i}$ are the second Love number and the compactness of the $i$-th object, respectively. The reduced tidal parameter $\tilde{\Lambda}$ is
$\tilde{\Lambda}=\frac{16}{13}\left[\frac{(m_{1}+12m_{2})m_{1}^{4}\Lambda_{1}}{M^{5}}+(1\leftrightarrow 2)\right]\,.$ (5)
The quadrupolar tidal polarizability $\kappa^{\rm T}_{2}$ is
$\kappa^{\rm T}_{2}=3\nu\,\left[\left(\frac{m_{1}}{M}\right)^{3}\Lambda_{1}+(1\leftrightarrow 2)\right]\,.$ (6)
The extrinsic parameters of the source ${\boldsymbol{\theta}_{\rm ext}}=\{D_{L},\iota,\alpha,\delta,\psi,t_{\rm mrg},\phi_{\rm mrg}\}$, i.e. inclination angle $\iota$, right ascension angle $\alpha$, declination angle $\delta$, polarization angle $\psi$, time of coalescence $t_{\rm mrg}$, and phase at the merger $\phi_{\rm mrg}$, allow us to identify the location and orientation of the source. Moreover, we include the recalibration parameters $\boldsymbol{\delta}_{\rm fit}$ that account for deviations from the predictions of the EoS-insensitive relations consistently with the related theoretical uncertainties. NRPMw is parameterized by two additional sets of degrees of freedom (see Paper I for a detailed discussion). The PM parameters ${\boldsymbol{\theta}_{\rm PM}^{\rm free}}=\{\phi_{\rm PM},t_{c},\alpha_{\rm drift}\}$ correspond to: the PM phase-shift $\phi_{\rm PM}$ that identifies the phase discontinuity after merger; the time of collapse $t_{c}$ that characterizes the collapse of the remnant into BH; and the frequency drift $\alpha_{\rm drift}$ that accounts for linear evolution of the dominant $f_{2}$ component.
II Method
II.1 Waveform model
Figure 1: Attachment between TEOBResumSPA and NRPMw for a GW template of non-spinning binary with $m_{1}=1.5~{}{\rm M_{\odot}}$, $m_{2}=1.4~{}{\rm M_{\odot}}$, $\Lambda_{1}=400$, $\Lambda_{2}=600$ and $D_{L}=40~{}{\rm Mpc}$. Since the model is defined in the frequency domain, we obtain $h_{22}(t)$ by performing an inverse FFT of $h_{22}(f)$.
In Paper I we introduced the PM NRPMw waveform, which can be employed to extend in the kiloHertz regime the existing frequency-domain approximants for GW signals from BNS mergers. For our studies, we resort to the EOB model TEOBResumS Nagar et al. (2018); Akcay et al. (2019); Nagar et al. (2019), in its frequency domain formulation, TEOBResumSPA Nagar and Rettegno (2019); Gamba et al. (2021a), for the representation of the IM part and present here the full spectrum template TEOBResumSPA_NRPMw.
We attach the EOB model for the IM to the PM NRPMw template, focusing on the dominant $(2,2)$ mode. The information required to perform the attachment of NRPMw with TEOBResumSPA is the time and the phase at merger, respectively $t_{\rm mrg}$ and $\phi_{\rm mrg}$. These data can be extracted from the EOB IM waveform. Then, resorting linear properties of the Fourier transform, the inspiral-merger-postmerger (IMPM) waveform can be computed as the sum of the pre-merger template $h_{\rm EOB}$ and the PM one $h_{\rm PM}$ in the frequency-domain,
$h_{\rm IMPM}(f)=h_{\rm EOB}(f)+h_{\rm PM}(f)\,{\rm e}^{-2\pi{\rm i}ft_{\rm mrg}+\phi_{\rm mrg}}\,,$ (7)
where $h_{\rm PM}(f)$ is taken such that its reference time and phase are vanishing.
Figure 1 illustrates an example of such an attachment procedure for a representative configuration with $M=2.9{\rm M_{\odot}}$, $q=1.072$, $\Lambda_{1}=400$ and $\Lambda_{2}=600$. As shown, the TEOBResumSPA_NRPMw waveform (dashed line) accurately reproduces the IM dynamics as described by the TEOBResumSPA template, as well as the PM segment following NRPMw. The attachment occurs after the merger, with the adoption of the phase from the EOB model, according to Eq. (7). To guarantee a smooth transition, the fusion wavelet, i.e. the portion of the NRPMw waveform between the merger time (as predicted by TEOBResumSPA) and $t_{0}$, defined as the first amplitude’s minimum after the merger (See Paper I), is used to perform a continuous attachment of the PM with the IM. Consequently, this segment is not compatible with either model.
We also validate the model against six NR simulations that are not included in the calibration set of NRPMw, finding that the model is $89-95\%$ faithful to NR when considering the full spectrum of frequencies and the ET-D sensitivity curve Hild et al. (2011); Hild (2012). More details on the IMPM model and its validation can be found in Appendix A.
II.2 Inference
Figure 2: Amplitude spectral density (ASD) of ET in configuration D Hild et al. (2011); Hild (2012) (in black) and comparison with ASD of LIGO L1 detector, relative to GW170817 event Abbott et al. (2019a) (in grey), to show the improvement in the detection of GW signals with XG detectors. The solid colored lines represent five exemplary BNS waveform spectra with TEOBResumSPA_NRPMw template for $\mathcal{M}=1.24~{}{\rm M_{\odot}}$, $q=1$, varying $\tilde{\Lambda}$ and $D_{L}$ as indicated in the legend.
We perform Bayesian parameter estimation (PE) experiments by injecting and analyzing mock zero-noise BNS signals with XG detectors. We employ the power spectral density (PSD) curve of the ET detector, as reported in Fig. 2, in configuration D Hild et al. (2011); Hild (2012), i.e. an equilateral triangular shape with $10~{}{\rm km}$ long arms. We simulate both IM-only and IMPM waveforms and recover them using the same model employed for the injection, TEOBResumSPA and TEOBResumSPA_NRPMw respectively.
Bayesian inference is performed with the parallel bajes pipeline Breschi et al. (2021). In order to efficiently explore the parameter space and compute the evidence, we utilize nested sampling algorithm Veitch et al. (2015); Handley et al. (2015); Allison and Dunkley (2014) as implemented in the dynesty library Speagle (2020) with a minimum of 2048 live points.
Bayesian analysis is based on the evaluation of the likelihood function (See Appendix B.1 for additional details), which is the most computationally expensive part. To speed up the computations, we make use of the relative binning technique Dai et al. (2018); Zackay et al. (2018) in the IM part. Relative binning consists in the precomputation of frequency-binned overlap of a fiducial waveform with the data, since the ratio of neighboring waveforms in frequency domain is smooth in the inspiral part. In particular, for the IM-only PEs we employ the relative binning method for the whole frequency range of the integration. For the IMPM cases, we utilize relative binning only up to a cutoff frequency $f_{\rm cut}=2048~{}{\rm Hz}$ and we perform a normal Bayesian PE without speed up after the cutoff frequency. The two likelihoods are then combined by summing their respective contributions:
$\log p({\text{{d}}}|\boldsymbol{\theta},H_{\rm S})=\log p({\text{{d}}}_{\rm IM}|\boldsymbol{\theta},H_{\rm IM}^{\rm RB})+\log p({\text{{d}}}_{\rm PM}|\boldsymbol{\theta},H_{\rm PM}).$ (8)
where $\log p({\text{{d}}}_{\rm IM}|\boldsymbol{\theta},H_{\rm IM}^{\rm RB})$ is the likelihood evaluated in the IM portion of the spectrum with relative binning, while $\log p({\text{{d}}}_{\rm PM}|\boldsymbol{\theta},H_{\rm PM})$ is evaluated in the high-frequency section, i.e. with NRPMw. For further explanation of the relative binning implementation and a discussion of the approximations used, we refer to Appendix B.2.
II.3 EoS constraints from quasi-universal relations
Given an estimation of the intrinsic BNS parameters, we show how XG detectors can enable precise constraints on the EoS of NSs. Inference of IM delivers posteriors of the masses of the compact objects and of the tidal polarizability parameters. The information contained in the IMPM signals, instead, allows to go beyond $m_{i}$ and $\Lambda_{i}$ and probe high-density NS properties. Breschi et al. (2022a) combined the Fisher matrix approach in the IM part with the Bayesian analysis of the PM in order to obtain a tighter constraint on the mass-radius relation. Here we perform a full spectrum Bayesian PE, aiming to obtain more reliable inference, particularly in the high-density regime. The resulting pipeline constrains different quantities of interest in EoS modeling and is described in what follows.
The first step is to map the posteriors for the total mass and the peak PM frequency $f_{2}$ into posteriors of maximum density ${\rho_{\rm max}^{\rm TOV}}$ and radius at the maximum mass ${R_{\rm max}^{\rm TOV}}$, using the NR phenomenological relations Breschi et al. (2022a, 2024b):
${\rho_{\rm max}^{\rm TOV}}=\frac{a_{0}c^{6}}{G^{3}M^{2}}\left[1+a_{1}\left(\frac{c^{3}}{\pi GMf_{2}}\right)^{1/6}\right]\,,$ (9)
with $(a_{0},a_{1})=(0.135905,-0.59506)$ and
$\begin{split}{R_{\rm max}^{\rm TOV}}/M&=(5.81\pm 0.13)-(123.4\pm 7.2)Mf_{2}\\ &+(1121\pm 99)(Mf_{2})^{2}\,.\end{split}$ (10)
Subsequently, we use a sample of two millions EoSs from Godzieba et al. (2021) generated under minimal assumptions, i.e. general relativity, causality, the maximum NS mass above the maximum NS mass from pulsar observations, ${M_{\rm max}^{\rm TOV}}>1.97{\rm M_{\odot}}$ Antoniadis et al. (2013), and consistent with the upper bound on the reduced tidal parameter $\tilde{\Lambda}$ from GW170817 Abbott et al. (2017a, 2019b, 2018b). We reweigh the EoS sample with the posteriors of the masses, the reduced tidal parameter, ${\rho_{\rm max}^{\rm TOV}}$ and ${R_{\rm max}^{\rm TOV}}$, employing emcee sampler Foreman-Mackey et al. (2013). This allows us to place constraints not only on the NS mass-radius diagram, but also to determine the 90% credibility region for the tidal deformability curve $\Lambda(M)$ and on the pressure-density relation $p(\rho)$. Specifically, from the EoS sampling we obtain a distribution of EoSs that most closely resemble the one used for the injection. For each sampled EoS, we compute the relations between the mass and the radius of the star and between the tidal polarizability parameter and the mass, yielding a 90% confidence interval (CI) on the $M(R)$ and the $\Lambda(M)$ curve, respectively. Using the same EoS distribution we solve the Tolmann-Oppenheimer-Volkoff (TOV) equations to determine the corresponding pressure-density relation, from which we derive the 90% confidence region for the $p(\rho)$ curve.
The maximum mass of non-rotating NSs, ${M_{\rm max}^{\rm TOV}}$, is a crucial parameter for constraining the mass-radius diagram and consequently the EoS. Additionally, it is also essential for the determination of the probability of prompt collapse. Specifically, PC is expected to occur when the total mass of the system exceeds a threshold mass, which is related through phenomenological relations to ${M_{\rm max}^{\rm TOV}}$ Hotokezaka et al. (2011); Bauswein et al. (2013):
$M_{\rm thr}=k_{\rm thr}(C_{\rm max}){M_{\rm max}^{\rm TOV}},$ (11)
where $k_{\rm thr}(C_{\rm max})$ is a function of the compactness of the star. The values of $C_{\rm max}$ and ${M_{\rm max}^{\rm TOV}}$ come from the EoS sampling as presented above. In our analysis we adopt the updated fits for $k_{\rm thr}$ of Kashyap et al. (2022), which were obtained for equal-mass binaries. Perego et al. (2022) provide a more general characterization with mass ratio effects and show that the latter give corrections compatible with the fit uncertainty down to mass ratios $q\lesssim 1.4-1.5$ (depending on EoS). For larger mass ratio PC is driven by accretion effects of the secondary stars to the primary and require a more sophisticated parametrization. Empirical models for the threshold mass can be obtained also in those cases by considering, for example, the EoS incompressibility parameter at extreme densities. Typical deviations from the equal-mass relation are up to ${\sim}20\%$ for mass ratios $q\sim 2$.
In our PC analyses, we consider equal-mass binaries and determine the $90\%$ CI of the maximum mass ${M_{\rm max}^{\rm TOV}}$ and the corresponding radius ${R_{\rm max}^{\rm TOV}}$ of the NS from the mass-radius diagram and insert them in Eq. (11). We then calculate the probability of PC as
$P_{\rm PC}=P(M>M_{\rm thr}|d)\,,$ (12)
as proposed by Agathos et al. (2020).
In addition, the presence of PC can also be statistically evaluated by computing the Bayes’ factor (BF) between models that assume either a PC to a BH or the formation of a NS remnant Breschi et al. (2019) (see Appendix B.1 for details on BF calculation). An alternative method to assess the occurrence of PC is to estimate the collapse time, $t_{c}$, to a BH, which is a parameter introduced in NRPMw waveforms. For practical purposes we assume that a system promptly collapses to a BH if $t_{c}<t_{c}^{\rm thr}=2~{}{\rm ms}$. This threshold just serves as an upper limit for the PC timescale; since a proper definition of PC requires inspection of the bulk merger dynamics Radice et al. (2018).
III Results
III.1 Inspiral-merger
Figure 3: Relative precision on chirp mass (top) and mass ratio (bottom) with respect to the different initial frequencies for a non-spinning BNS with $\mathcal{M}=1.1976917{\rm M_{\odot}}$, $q=1.5$, $\tilde{\Lambda}=488$, at different luminosity distances. A lower initial frequency or a closer source leads to improved precision in both the chirp mass and the mass ratio estimations.
Figure 4: Posterior distributions for reduced tidal parameter $\tilde{\Lambda}$ with respect to SNR, varying the injected luminosity distances, for eight fiducial binaries. The initial frequency is kept fixed to $f_{0}=5~{}{\rm Hz}$. As the SNR increases, i.e. the luminosity distance decreases, the width of the distribution of $\tilde{\Lambda}$ shrinks.
Figure 5: Posterior distributions for reduced tidal parameter $\tilde{\Lambda}$ with respect to SNR, varying the initial frequencies of the detection, for eight fiducial binaries. The luminosity distance is kept fixed to $D_{L}=40~{}{\rm Mpc}$. As the SNR increases, i.e. the initial frequency decreases, the width of the distribution of $\tilde{\Lambda}$ shrinks. Two different configurations of ET, triangular Hild et al. (2011); Hild (2012) in orange and two-L Branchesi et al. (2023) in blue, give comparable results.
Figure 6: Comparison of the posterior distributions for extrinsic parameters for two different ET configurations: triangular Hild et al. (2011); Hild (2012) in orange and two-L Branchesi et al. (2023) in blue. The injections have been performed with TEOBResumSPA template with $\mathcal{M}=1.344787{\rm M_{\odot}}$, $q=1.21$, $\tilde{\Lambda}=172$, $D_{L}=40\,{\rm Mpc}$ and $\iota=45^{\circ}$. The initial frequency $f_{0}$ is $10\,{\rm Hz}$. We are showing the differences between the recovered parameters and the injected ones, since the sky position corresponds to the relative best location for each detector. The posterior distributions are comparable, except for the estimation of the declination, where the two-L configuration provides tighter constraints.
Table 1: Summary of the PEs with TEOBResumSPA for a BNS with $\mathcal{M}=1.1976917{\rm M_{\odot}}$, $q=1.5$, $\tilde{\Lambda}=488$, $\delta\tilde{\Lambda}=-60$, $\chi_{\rm eff}=0$, $\iota=0$. The first three columns report the injected luminosity distance, the initial frequency of the analyses and the SNR of the injected data; the last seven columns report the recovered values, with the median of the posterior distributions and the $90\%$ CIs.
Injected properties Recovered values
$D_{L}~{}[{\rm Mpc}]$ $f_{0}{[\rm Hz]}$ SNR $\mathcal{M}~{}[{\rm M_{\odot}}]$ $q$ $\tilde{\Lambda}$ $\delta\tilde{\Lambda}$ $\chi_{\rm eff}$ $D_{L}~{}[{\rm Mpc}]$ $\iota$
2 2261 $1.1976917^{+0.0000002}_{-0.0000002}$ $1.500^{+0.002}_{-0.002}$ $490^{+9}_{-9}$ $-50^{+30}_{-30}$ $-0.0001^{+0.0001}_{-0.0001}$ $38.5^{+1.0}_{-0.8}$ $0.26^{+0.07}_{-0.13}$
40 5 2228 $1.1976917^{+0.0000005}_{-0.0000005}$ $1.500^{+0.004}_{-0.004}$ $488^{+10}_{-10}$ $-70^{+20}_{-20}$ $0.0000^{+0.0002}_{-0.0002}$ $40.00^{+0.01}_{-0.01}$ $0.02^{+0.01}_{-0.01}$
10 1613 $1.197691^{+0.000002}_{-0.000002}$ $1.498^{+0.011}_{-0.009}$ $490^{+12}_{-14}$ $-60^{+30}_{-30}$ $-0.0001^{+0.0009}_{-0.0007}$ $40.01^{+0.02}_{-0.01}$ $0.02^{+0.01}_{-0.01}$
20 1301 $1.19769^{+0.00001}_{-0.00001}$ $1.50^{+0.04}_{-0.05}$ $490^{+30}_{-20}$ $-60^{+70}_{-30}$ $-0.0002^{+0.0030}_{-0.0040}$ $37.7^{+0.9}_{-0.8}$ $0.33^{+0.06}_{-0.08}$
2 1130 $1.1976917^{+0.0000004}_{-0.0000004}$ $1.500^{+0.003}_{-0.004}$ $490^{+20}_{-20}$ $-40^{+65}_{-40}$ $-0.0001^{+0.0002}_{-0.0003}$ $76^{+2}_{-3}$ $0.30^{+0.09}_{-0.13}$
80 5 1114 $1.1976915^{+0.0000010}_{-0.0000010}$ $1.499^{+0.008}_{-0.007}$ $490^{+30}_{-20}$ $-40^{+80}_{-50}$ $-0.0001^{+0.0005}_{-0.0005}$ $76^{+3}_{-2}$ $0.31^{+0.09}_{-0.14}$
10 806 $1.197690^{+0.000005}_{-0.000006}$ $1.49^{+0.03}_{-0.03}$ $500^{+50}_{-30}$ $-40^{160}_{-50}$ $-0.0005^{+0.0020}_{-0.0020}$ $75^{+3}_{-3}$ $0.33^{+0.09}_{-0.14}$
20 650 $1.19769^{+0.00002}_{-0.00003}$ $1.49^{+0.06}_{-0.14}$ $500^{+90}_{-40}$ $-30^{+220}_{-60}$ $-0.0007^{+0.0060}_{-0.0110}$ $75^{+3.0}_{-3}$ $0.36^{+0.09}_{-0.13}$
2 754 $1.1976916^{+0.0000006}_{-0.0000006}$ $1.498^{+0.006}_{-0.006}$ $500^{+30}_{-30}$ $0^{+110}_{-90}$ $-0.0000^{+0.0007}_{-0.0006}$ $113^{+5}_{-4}$ $0.3^{+0.1}_{-0.2}$
120 5 743 $1.197692^{+0.000001}_{-0.000002}$ $1.50^{+0.01}_{-0.01}$ $490^{+40}_{-30}$ $-30^{+150}_{-70}$ $-0.0001^{+0.0008}_{-0.0010}$ $113^{+4}_{-5}$ $0.3^{+0.1}_{-0.1}$
10 538 $1.197689^{+0.000008}_{-0.000008}$ $1.49^{+0.04}_{-0.04}$ $510^{+60}_{-50}$ $0^{+170}_{-100}$ $-0.001^{+0.003}_{-0.003}$ $112^{+6}_{-5}$ $0.4^{+0.1}_{-0.2}$
20 434 $1.19768^{+0.00002}_{-0.00004}$ $1.46^{+0.09}_{-0.14}$ $530^{+70}_{-70}$ $20^{+200}_{-110}$ $-0.003^{+0.008}_{-0.011}$ $111^{+6}_{-5}$ $0.4^{+0.1}_{-0.2}$
2 301 $1.197692^{+0.000002}_{-0.000002}$ $1.50^{+0.01}_{-0.01}$ $510^{+60}_{-50}$ $80^{+160}_{-160}$ $-0.0004^{+0.0010}_{-0.0013}$ $270^{+10}_{-10}$ $0.4^{+0.1}_{-0.2}$
300 5 297 $1.197691^{+0.000003}_{-0.000004}$ $1.49^{+0.02}_{-0.03}$ $530^{+50}_{-70}$ $100^{130}_{-180}$ $-0.0002^{+0.0020}_{-0.0020}$ $280^{+20}_{-20}$ $0.4^{+0.2}_{-0.2}$
10 215 $1.19769^{+0.00001}_{-0.00002}$ $1.49^{+0.07}_{-0.11}$ $500^{+100}_{-70}$ $20^{+210}_{-130}$ $-0.0005^{+0.0050}_{-0.0090}$ $270^{+20}_{-20}$ $0.4^{+0.1}_{-0.3}$
20 173 $1.19768^{+0.00008}_{-0.00007}$ $1.4^{+0.2}_{-0.3}$ $540^{+70}_{-110}$ $100^{+120}_{-220}$ $-0.005^{+0.021}_{-0.020}$ $270^{+20}_{-20}$ $0.4^{+0.2}_{-0.2}$
In this section we consider nine injection-recovery PEs with TEOBResumSPA. We vary the initial frequency of the analyses and the luminosity distance of the binaries, comparing the recovered parameters with the injected ones. This approach allows us to investigate how constraints on various parameters depend on the different SNRs of the signals.
We generate artificial data in the frequency domain, using the triangular, triple-interferometer ET detector Hild et al. (2008); Punturo et al. (2010b); Maggiore et al. (2020), with ET-D sensitivity Hild et al. (2011); Hild (2012) and a sampling rate of $4096~{}{\rm Hz}$. The sky location corresponds to the optimal location for the ET detector $\{\alpha=2.64,\delta=0.762\}$ and the binaries are oriented with $\iota=0$ and $\psi=0$. The injections are performed in zero-noise, including only the dominant (2,2) mode. The injected strain is analyzed in the frequency domain from $f_{0}$ to $2048~{}{\rm Hz}$, where $f_{0}\in[2,5,10,20]~{}{\rm Hz}$. We set the priors for chirp mass $\mathcal{M}\in[0.8,1.5]{\rm M_{\odot}}$, for mass ratio $q\in[1,2]$ and for the tidal polarizability parameters $\Lambda_{1,2}\in[0,3000]$. The spins are aligned with the orbital angular momentum, with an isotropic prior $\chi_{1,2}\in[-0.05,0.05]$ for the $z-$component. Finally, for the luminosity distance we employ volumetric priors $D_{L}\in[10,150]~{}{\rm Mpc}$, except when the injected luminosity distance is $300~{}{\rm Mpc}$, where we use $D_{L}\in[200,400]~{}{\rm Mpc}$. We marginalize over reference phase.
III.1.1 Source parameters
Figure 3 shows the relative precision on chirp mass and mass ratio calculated as
$\frac{\Delta x}{x}=\frac{x_{95\%}-x_{5\%}}{x_{m}},$ (13)
where $x_{m}$ is the median value and $x_{95\%}$ ($x_{5\%}$) corresponds to the $95$th ($5$th) percentile of the posterior distribution of the parameter $x$. The injected GW signals correspond to non-spinning BNS with $\mathcal{M}=1.1976917{\rm M_{\odot}}$, $q=1.5$ and $\tilde{\Lambda}=488$. We vary the luminosity distance in $[40,80,120,300]~{}{\rm Mpc}$. The precision on both $\mathcal{M}$ and $q$ increases as the distance of the source decreases, since the SNR is higher. Similarly, keeping a fixed luminosity distance, the precision increases as the initial frequency $f_{0}$ of the analysis decreases, e.g. at $D_{L}=300~{}{\rm Mpc}$, decreasing the frequency from $20~{}{\rm Hz}$ to $5~{}{\rm Hz}$, the precision on the measurement of both chirp mass and mass ratio improves of about one order of magnitude.
Figure 4 shows the distributions of the differences between the recovered and the injected values of the reduced tidal parameter,
$\frac{\Delta\tilde{\Lambda}}{\tilde{\Lambda}}=\frac{|\tilde{\Lambda}_{\rm rec}-\tilde{\Lambda}_{\rm inj}|}{\tilde{\Lambda}_{\rm inj}}\,,$ (14)
with respect to the SNR, for eight different binaries. As the SNR increases, that is when the luminosity distance decreases, the width of the distribution shrinks, meaning that the measurement of $\tilde{\Lambda}$ is more precise. This is mostly due to the improvement in the chirp mass and mass ratio discussed above. We run the same test fixing the luminosity distance and varying the initial frequencies of the analyses. The results are presented in Fig. 5. We obtain a similar result as before: as the SNR increases, namely as the initial frequency decreases, it is possible to have a better estimation on the value of $\tilde{\Lambda}$, e.g. of about a factor two, when the initial frequency changes from 20 to $5~{}{\rm Hz}$. Reducing the luminosity distance from $120~{}{\rm Mpc}$ to $40~{}{\rm Mpc}$, instead, improves the precision on the reduced tidal parameters by a factor between two and three. At the same time, we notice that having a larger luminosity distance, i.e. 80 Mpc, but a smaller initial frequency, i.e. 5 Hz, enhances the precision in the measurement of all the parameters with respect to a closer source ($D_{L}=40~{}{\rm Mpc}$) at $f_{0}=20~{}{\rm Hz}$, even if the SNR is higher in the latter case. Thus, decreasing the initial frequency is an effective strategy to achieve better constraints in the estimation of the binary parameters.
Table 1 presents the recovered median values with the $90\%$ confidence regions for the inferred parameters. In addition to the consideration above about the chirp mass and the mass ratio, as the initial frequency decreases from 20 Hz to 2 Hz, the precision on the estimation of the reduced tidal parameter increases of about a factor two and the precision on the effective spin of one order of magnitude. The precision of the luminosity distance and of the inclination, instead, is comparable for the different cases and the inferred values are not always compatible with the injected ones. This is due to the known degeneracy between distance and inclination, which arises when considering waveforms constructed only with the $(2,\pm 2)$ modes Cutler and Flanagan (1994). Despite this, our results of the precision of the luminosity distance are consistent with those reported in Singh and Bulik (2022).
In Table 1 and in Table LABEL:tab:IM2 of Appendix C we also report the recovered values of the asymmetric tidal parameter $\delta\tilde{\Lambda}$. The estimation on this parameter carries substantial uncertainties, in some case larger than 100%. The trend in the improvement of its precision follows the same behavior discussed above for the reduced tidal parameter. We notice that, when the SNR is sufficiently high, i.e. SNR1500greater-than-or-equivalent-toabsent1500$\gtrsim$ 1500≳ 1500 for $q>1.2$, the recovered values of $\delta\tilde{\Lambda}$ deviate significantly from zero, indicating a measurable asymmetry in the tidal polarizability parameters. Furthermore, the joint estimation of $\delta\tilde{\Lambda}$ and $\tilde{\Lambda}$ leads to constraints on the individual tidal polarizability parameters. In few cases, when the mass ratio is $q\gtrsim 1.289$ and the SNR has the highest value, i.e. $f_{0}=$ 5Hz and $D_{L}=$ 40Mpc, we notice that both the posterior distributions of $\Lambda_{1}$ and $\Lambda_{2}$ are not compatible with zero, enabling us to confidently identify the source as a BNS merger. However, in all the other cases, the posterior distribution of $\Lambda_{1}$ remains compatible with zero, preventing us from discarding the BH-NS hypothesis.
III.1.2 Comparison of ET configurations
In Figure 5 we compare two different configurations of ET, the triangular one Hild et al. (2011); Hild (2012), described in Sec. II.2 and a two-L configuration Branchesi et al. (2023), where the first detector is located in the Meuse-Rhine region and the second one in Sardinia, with the same sensitivity as the triangular one Hild et al. (2011); Hild (2012) and a relative orientation between the two detectors of $45^{\circ}$. We find that the posterior distributions of $\tilde{\Lambda}$ for the two configurations are comparable. Our results are in line with Puecher et al. (2023b).
We then perform a comparison between the two different configurations of ET for what concerns the recovery of the extrinsic parameters, i.e. right ascension $\alpha$, declination $\delta$, inclination $\iota$ and luminosity distance $D_{L}$. Figure 6 shows the posterior distributions of the difference between the recovered and the injected values for the extrinsic parameters, since for each detector we use the best location. We notice that the $90\%$ confidence regions are comparable for the two configurations for all the parameters, except for the declination, which is better recovered in the case of the two-L configuration.
Branchesi et al. (2023); Begnoni et al. (2025) and Puecher et al. (2023b) compared various ET geometries using a Fisher matrix approach in the former cases and a Bayesian PE in the latter. They all concluded that an improvement in the precision on the inference of the source parameters comes primarily from the arms length, while changing shape or alignment of the detectors plays a minimal role in the accuracy increase. Additionally, Branchesi et al. (2023) found that the estimation of the source parameters is comparable for the two configurations we considered, while the triangular performs better in the estimation of the inclination angle. Regarding the angular localization, their results indicate that the two-L is better compared to the triangular one. This improvement is attributed to the ability of the two-L setup to partially triangulate, thanks to its longer baseline. These findings are compatible with our results.
III.2 Inspiral-merger-postmerger
Table 2: Summary of the PEs with TEOBResumSPA for IM and TEOBResumSPA_NRPMw for IMPM case. The first five columns report the injected parameters, included the EoS used; the central six columns report the recovered values, with the median of the posterior distributions and the $90\%$ CIs and the last column displays the BFs, with statistical uncertainties as in Skilling (2006).
Injected properties Recovered values IM Recovered values IMPM
morph. $\mathcal{M}~{}[{\rm M_{\odot}}]$ $q$ $\tilde{\Lambda}$ SNR $\mathcal{M}~{}[{\rm M_{\odot}}]$ $q$ $\tilde{\Lambda}$ $\mathcal{M}~{}[{\rm M_{\odot}}]$ $q$ $\tilde{\Lambda}$ $\log\mathcal{B}_{\rm{IM}}^{\rm{IMPM}}$
DD2 1.339520 1 403 1427 $1.339519^{+0.000001}_{-0.000001}$ $1.013^{+0.015}_{-0.012}$ $400^{+4}_{-6}$ $1.339519^{+0.000001}_{-0.000001}$ $1.016^{+0.012}_{-0.011}$ $400^{+4}_{-6}$ $32.29^{+0.47}_{-0.47}$
[…]
DD2 1.522426 1 178 1583 $1.522425^{+0.000001}_{-0.000001}$ $1.011^{+0.013}_{-0.010}$ $177^{+2}_{-3}$ $1.522425^{+0.000001}_{-0.000002}$ $1.016^{+0.012}_{-0.011}$ $176^{+3}_{-3}$ $29.54^{+0.48}_{-0.48}$
LS220 1.170202 1 703 1278 $1.170202^{+0.000001}_{-0.000001}$ $1.015^{+0.014}_{-0.013}$ $699^{+6}_{-8}$ $1.170202^{+0.000001}_{-0.000001}$ $1.015^{+0.012}_{-0.012}$ $699^{+7}_{-8}$ $24.99^{+0.47}_{-0.47}$
DD2 1.315304 1.289 442 1405 $1.315304^{+0.000002}_{-0.000002}$ $1.289^{+0.002}_{-0.002}$ $443^{+12}_{-14}$ $1.315304^{+0.000002}_{-0.000002}$ $1.290^{+0.002}_{-0.002}$ $440^{+12}_{-13}$ $22.25^{+0.48}_{-0.48}$
SLy 1.175243 1 394 1466 $1.175243^{+0.000001}_{-0.000001}$ $1.014^{+0.014}_{-0.012}$ $390^{+5}_{-6}$ $1.175243^{+0.000001}_{-0.000001}$ $1.014^{+0.011}_{-0.011}$ $392^{+4}_{-5}$ $19.39^{+0.48}_{-0.48}$
BLh 1.227164 1 412 1329 $1.227163^{+0.000001}_{-0.000001}$ $1.012^{+0.015}_{-0.011}$ $409^{+5}_{-6}$ $1.227163^{+0.000001}_{-0.000001}$ $1.017^{+0.012}_{-0.012}$ $408^{+4}_{-6}$ $18.90^{+0.47}_{-0.47}$
DD2 1.043690 1 1638 1164 $1.043690^{+0.000000}_{-0.000001}$ $1.015^{+0.016}_{-0.013}$ $1631^{+9}_{-18}$ $1.043690^{+0.000001}_{-0.000001}$ $1.018^{+0.014}_{-0.013}$ $1629^{+10}_{-17}$ $16.72^{+0.46}_{-0.46}$
BLh 1.341381 1 222 1429 $1.341381^{+0.000001}_{-0.000001}$ $1.012^{+0.013}_{-0.011}$ $220^{+3}_{-4}$ $1.341380^{+0.000001}_{-0.000001}$ $1.013^{+0.012}_{-0.010}$ $221^{+3}_{-4}$ $15.16^{+0.48}_{-0.48}$
SFHo 1.130733 1 538 1243 $1.130733^{+0.000000}_{-0.000001}$ $1.014^{+0.015}_{-0.012}$ $534^{+6}_{-8}$ $1.130732^{+0.000001}_{-0.000001}$ $1.017^{+0.012}_{-0.012}$ $534^{+6}_{-8}$ $15.05^{+0.46}_{-0.46}$
BLh 1.239566 1 385 1339 $1.239565^{+0.000001}_{-0.000001}$ $1.117^{+0.004}_{-0.004}$ $385^{+7}_{-12}$ $1.239565^{+0.000001}_{-0.000001}$ $1.117^{+0.004}_{-0.003}$ $385^{+6}_{-9}$ $14.36^{+0.47}_{-0.47}$
SFHo 1.273413 1 249 1369 $1.273412^{+0.000001}_{-0.000001}$ $1.013^{+0.014}_{-0.011}$ $247^{+3}_{-4}$ $1.273412^{+0.000001}_{-0.000001}$ $1.016^{+0.013}_{-0.012}$ $246^{+4}_{-5}$ $14.19^{+0.47}_{-0.47}$
SFHo 1.344787 1.212 172 1432 $1.344787^{+0.000002}_{-0.000002}$ $1.213^{+0.002}_{-0.002}$ $171^{+6}_{-9}$ $1.344787^{+0.000002}_{-0.000001}$ $1.213^{+0.002}_{-0.002}$ $171^{+5}_{-7}$ $13.94^{+0.48}_{-0.48}$
SLy 1.058974 1 743 1178 $1.058974^{+0.000000}_{-0.000001}$ $1.013^{+0.015}_{-0.012}$ $738^{+7}_{-10}$ $1.058974^{+0.000001}_{-0.000001}$ $1.018^{+0.012}_{-0.013}$ $737^{+7}_{-10}$ $12.99^{+0.46}_{-0.46}$
SLy 1.244295 1 264 1344 $1.244295^{+0.000001}_{-0.000001}$ $1.012^{+0.014}_{-0.011}$ $262^{+4}_{-4}$ $1.244295^{+0.000001}_{-0.000001}$ $1.016^{+0.013}_{-0.012}$ $261^{+4}_{-5}$ $12.89^{+0.47}_{-0.47}$
SLy 1.303371 1 191 1396 $1.303371^{+0.000001}_{-0.000001}$ $1.012^{+0.013}_{-0.011}$ $189^{+3}_{-4}$ $1.303370^{+0.000001}_{-0.000001}$ $1.015^{+0.012}_{-0.011}$ $189^{+3}_{-4}$ $12.51^{+0.47}_{-0.47}$
LS220 1.223692 1.287 525 1325 $1.223692^{+0.000001}_{-0.000002}$ $1.288^{+0.002}_{-0.002}$ $525^{+13}_{-16}$ $1.223692^{+0.000001}_{-0.000001}$ $1.288^{+0.002}_{-0.002}$ $525^{+11}_{-15}$ $12.22^{+0.47}_{-0.47}$
SLy 1.304020 1.222 192 1396 $1.304020^{+0.000001}_{-0.000002}$ $1.220^{+0.002}_{-0.002}$ $192^{+6}_{-9}$ $1.304020^{+0.000001}_{-0.000001}$ $1.220^{+0.002}_{-0.002}$ $192^{+5}_{-7}$ $9.78^{+0.48}_{-0.48}$
SLy 1.183075 1.257 365 1289 $1.183075^{+0.000001}_{-0.000001}$ $1.257^{+0.002}_{-0.002}$ $365^{+10}_{-14}$ $1.183075^{+0.000001}_{-0.000001}$ $1.257^{+0.002}_{-0.002}$ $363^{+11}_{-14}$ $8.91^{+0.47}_{-0.47}$
SFHo 1.339422 1 130 1427 $1.339422^{+0.000001}_{-0.000001}$ $1.012^{+0.013}_{-0.011}$ $129^{+3}_{-3}$ $1.339422^{+0.000001}_{-0.000001}$ $1.014^{+0.012}_{-0.011}$ $129^{+3}_{-3}$ $7.17^{+0.48}_{-0.48}$
2B 1.175243 1 126 1283 $1.175243^{+0.000001}_{-0.000001}$ $1.013^{+0.013}_{-0.011}$ $124^{+4}_{-4}$ $1.175243^{+0.000001}_{-0.000001}$ $1.013^{+0.013}_{-0.011}$ $124^{+4}_{-4}$ $-3.63^{+0.47}_{-0.47}$
DD2 1.522426 1 178 1583 $1.522425^{+0.000001}_{-0.000002}$ $1.015^{+0.013}_{-0.011}$ $177^{+2}_{-3}$ $1.522425^{+0.000001}_{-0.000002}$ $1.015^{+0.013}_{-0.011}$ $177^{+2}_{-3}$ $-5.84^{+0.47}_{-0.47}$
SLy 1.303371 1 191 1396 $1.303370^{+0.000001}_{-0.000001}$ $1.014^{+0.013}_{-0.011}$ $189^{+3}_{-4}$ $1.303370^{+0.000001}_{-0.000001}$ $1.014^{+0.013}_{-0.011}$ $189^{+3}_{-4}$ $-6.28^{+0.47}_{-0.47}$
Figure 7: Posterior distributions for intrinsic parameters recovered from the injection of a BNS using LS220 EoS ($\mathcal{M}=1.223692\,{\rm M_{\odot}}$), as in Table 2. The recovery is performed both with TEOBResumSPA (IM) and TEOBResumSPA_NRPMw (IMPM). The injected values are in black. The posterior distributions obtained with the two models are analogous, since most of the information is included in the IM section.
Figure 8: Injected and recovered spectra of five exemplary binaries, with different EoSs. The solid colored lines report the injected waveforms, the shadowed regions correspond to the $90\%$ credibility regions and the vertical black dotted lines represent the PM peaks at $f_{2}$. We also report the ASD of the ET detector Hild et al. (2011); Hild (2012). Both the inspiral part and the peak of the PM are well recovered in all the cases.
This section is dedicated to the analysis of IMPM injection-recovery PEs. We employed the complete TEOBResumSPA_NRPMw template for the injections of GWs from different binaries, i.e. with different EoSs. As in the IM-only case, the detector used is ET, with triangular shape and ET-D sensitivity Hild et al. (2011); Hild (2012). The sky location is the same as in III.1 and the luminosity distance is $40~{}{\rm Mpc}$. The sampling rate is $16384~{}{\rm Hz}$ and the analyses are performed in the frequency range $f\in[20,8192]~{}\rm{Hz}$. The priors for chirp mass are $\mathcal{M}\in[0.8,2]\,{\rm M_{\odot}}$, for mass ratio $q\in[1,4]$, for the tidal polarizability parameters $\Lambda_{1,2}\in[0,4000]$ and for the luminosity distance are volumetric $D_{L}\in[10,1000]~{}{\rm Mpc}$. We perform the inference under the assumption of non-spinning binaries and we keep all the other extrinsic parameters fixed to the injected values. We add the recalibration parameters $\boldsymbol{\delta}_{\rm fit}$ to account for intrinsic errors of the calibrated formulae (see Paper I).
Figure 7 displays an example of posterior distributions for the LS220 binary of Table 2 with $\mathcal{M}=1.223692\,{\rm M_{\odot}}$. We compare between IM and IMPM recoveries. The posteriors distributions are, in both cases, correctly centered on the injected values (black lines) and they are similar. We observe that the inclusion of the PM does not significantly improve the constraints on chirp mass, mass ratio and tidal polarizability parameter compared to the IM-only PE. This outcome is primarily due to the dominant contribution of the IM segment, which exhibits a SNR on the order of $10^{3}$, whereas the SNR of the PM is only on the order of ten (the SNR threshold for detectability of the PM is 7 [Paper II]). Nevertheless, having a model that incorporates the PM remains essential for distinguishing between the characteristic PM features of a NS remnant and the case of a PC, by using BFs for model selection, as discussed below.
Figure 8 presents five illustrative recovered spectra, in which the injected waveforms are included within the $90\%$ CIs of the reconstructed spectra. Particularly, both the inspiral part and the frequency of the peak of the PM, $f_{2}$, are well recovered with tight error bars for all the binaries, i.e. it is possible to obtain constraints on $f_{2}$ for different EoSs with an uncertainty smaller than 8%. The high-frequency features and the region between the merger and the peak of the PM are consistent with the injection, but, in these cases, the $90\%$ CI are broader, indicating lower precision on the detection compared to the inspiral segment. This reduced precision arises because the strain amplitude is about two orders of magnitude smaller and the detector sensitivity is lower in the high-frequency region.
In Table 2 we report the results of the PE analyses we performed. In all the cases we use both the IM waveform TEOBResumSPA and the IMPM waveform TEOBResumSPA_NRPMw for the recoveries. The inferred values for the two cases are comparable and always compatible with the injected values. We compute the BFs as
$\log\mathcal{B}_{\rm{IM}}^{\rm{IMPM}}=\log\frac{p(d|{\rm IMPM})}{p(d|{\rm IM})}\,,$ (15)
where $p(d|H)$ is the evidence. In the upper part of the Table the BFs are larger than zero, meaning that the IMPM model is preferred over the IM model. All the BF values are on the order of ten, whereas the associated uncertainties, estimated with the criterion introduced in Skilling (2006), are on the order $10^{-1}$. This indicates a high degree of statistical confidence, supporting a clear detection of the presence of the PM. In the bottom part of the Table, instead, we test the ability of our pipeline to recover PC. The relative discussion is postponed to Sec. III.4.
III.3 EoS constraints
Figure 9: Posterior distributions of the difference between recovered and theoretical values for maximum density ${\rho_{\rm max}^{\rm TOV}}$ and radius at the maximum mass ${R_{\rm max}^{\rm TOV}}$ for five different EoSs. The recovered values are computed with Eqs. (9) and (10) and are consistent with the theoretical values from the EoS sequences, as reported in the legend.
[…]
In this section we present the results of the EoS study enabled by our PE analyses. In particular EoS sampling can be performed starting from the results of the IM-only PEs, employing only the masses and the tidal polarizability parameters, or from the results of the PM, incorporating also the constraints from the quasi-universal relations for the maximum density ${\rho_{\rm max}^{\rm TOV}}$ and the radius at the maximum mass ${R_{\rm max}^{\rm TOV}}$, as described in Sec. II.3.
We begin examining the distribution of the maximum density and the radius at the maximum mass, computed from the values of $f_{2}$ using the quasi-universal relations (9) and (10), respectively. Figure 9 displays the posterior distributions of $\Delta{\rho_{\rm max}^{\rm TOV}}=\rho_{\rm max}^{\rm TOV,PE}-{\rho_{\rm max}^{\rm TOV}}$ and $\Delta{R_{\rm max}^{\rm TOV}}=R_{\rm max}^{\rm TOV,PE}-{R_{\rm max}^{\rm TOV}}$, for five different EoSs. ${\rho_{\rm max}^{\rm TOV}}$ and ${R_{\rm max}^{\rm TOV}}$ corresponds to the maximal theoretical values obtained from the stable EoS sequence, while $\rho_{\rm max}^{\rm TOV,PE}$ and $R_{\rm max}^{\rm TOV,PE}$ are computed from the fits. We observe that for soft EoSs the distributions of $\Delta{\rho_{\rm max}^{\rm TOV}}$ are exactly centered around zero, whereas for stiffer EoSs, such as DD2, the distributions are shifted away from zero, although they remain compatible with zero within the $90\%$ CI. This discrepancy arises from the fit relations used in the analysis. In particular examining Fig. 1 of Breschi et al. (2022a), we note that the specific binary with DD2 EoS we used exhibits a larger deviation from the fitted relation for ${\rho_{\rm max}^{\rm TOV}}$, although it remains within the 90% CI.
Combining the posteriors of the total mass and the tidal parameters with the results above, we constrain the mass-radius diagram. Figure 10 shows the $M(R)$ relation for the SLy ($\mathcal{M}=1.244295~{}{\rm M_{\odot}}$) and BLh ($\mathcal{M}=1.227164~{}{\rm M_{\odot}}$) EoSs (See Table 2). The constrained areas are obtained with TEOBResumSPA for IM-only and TEOBResumSPA_NRPMw for IMPM. For PM-only we use the TEOBResumSPA_NRPMw PE to compute ${R_{\rm max}^{\rm TOV}}$ and ${\rho_{\rm max}^{\rm TOV}}$, but we are not including the constraints on the mass and the tidal polarizability parameter. We utilize the EoSs prior sample of Godzieba et al. (2021) and in our 90% CI 14372 EoSs are included. We additionally test scenarios varying this amount, i.e. by doubling it, and obtain consistent results.
The PM gives information at higher densities and together with the IM produces a tighter constraint around the theoretical mass-radius curve, giving in particular a better estimation of the maximum mass ${M_{\rm max}^{\rm TOV}}$ and the corresponding radius ${R_{\rm max}^{\rm TOV}}$. For IMPM, we find the CIs to be ${M_{\rm max}^{\rm TOV}}=2.03^{+0.07}_{-0.05}~{}{\rm M_{\odot}}$ and ${R_{\rm max}^{\rm TOV}}=10.23^{+0.25}_{-0.12}{\rm km}$ for SLy and ${M_{\rm max}^{\rm TOV}}=2.07^{+0.09}_{-0.08}~{}{\rm M_{\odot}}$ and ${R_{\rm max}^{\rm TOV}}=10.75^{+0.32}_{-0.25}~{}{\rm km}$ for BLh. The values we obtain are compatible with Breschi et al. (2021) and the relative error is roughly the same. Compared with earlier analyses based on GW170817 Breschi et al. (2024b); Nathanail et al. (2021); Koehn et al. (2025); Essick et al. (2020), an improvement by a factor three in the uncertainties of the maximum mass is achieved with XG detectors. For ${R_{\rm max}^{\rm TOV}}$, the precision enhances by a factor four Rutherford et al. (2024); Fan et al. (2024). We note that in the region of $1.54{\rm M_{\odot}}\lesssim M\lesssim 1.87{\rm M_{\odot}}$, the constrained IMPM 90% confidence region is not totally consistent with the theoretical mass-radius relation for SLy and BLh. We attribute this slight incompatibility to the sample of EoSs, that does not include EoSs exactly compatible with SLy or BLh at $M\lesssim 0.6~{}{\rm M_{\odot}}$.
Furthermore, Figure 11 reports the $\Lambda(M)$ curve derived from the combined IMPM analysis, evaluated for an equal-mass binary with BLh EoS, $M=2.8{\rm M_{\odot}}$ and $\tilde{\Lambda}=412$. The resulting constraint on the tidal polarizability parameter as a function of the mass is particularly tight around the theoretical curve for BLh, highlighting both the accuracy of our method in recovering the theoretical prediction and its robustness in probing the dense matter properties of NSs. From this analysis, we also extract constraints on the tidal polarizability parameter at $M=1.4{\rm M_{\odot}}$, finding that it lies within the range $529289<\Lambda_{1.4}<$ 529.
Finally, Figure 12 displays an example of direct inference of the pressure-density EoS relation. The theoretical SLy curve lies within the 90% CIs for both IM and IMPM cases. We observe a similar behavior to that seen in the mass-radius curve: the confidence region is narrower in the region with $5\times 10^{14}~{}{\rm g~{}cm^{-3}}\lesssim\rho\lesssim 13\times 10^{14}~{}{\rm g~{}cm^{-3}}$, while it becomes broader at both lower and higher densities. As in the other inferences discussed, the IMPM analysis improves the precision of the inference on $p(\rho)$ at higher densities.
III.4 Prompt collapse
Figure 13: Cumulative posterior distributions of the ratio between the total mass and the threshold mass for two different binaries. The probability of prompt collapse is represented by the fraction of posterior distribution that lies above the black line. In the case of SLy with $M=3~{}{\rm M_{\odot}}>M_{\rm thr}$ the probability of PC is larger than 80%, in the other case ($M=2.7~{}{\rm M_{\odot}}<M_{\rm thr}$) the probability is smaller than 35%. The difference between IM and IMPM comes from the different ability to constrain the maximum mass, between the two cases, as presented in Sec. III.3. The colored area represents the probability distribution for the promptly collapsing binary of having a collapse time larger than the threshold collapse time, $t_{\rm c}^{\rm thr}=2~{}{\rm ms}$. The whole area lies below the black line, therefore the probability is vanishing.
In this section we discuss the identification of PC. We initially analyze the lower part of Table 2, for which we injected GW signals with the IM-only model TEOBResumSPA, to simulate a PC. The recovery is performed both with IM and IMPM templates. In these cases the BFs $\log\mathcal{B}_{\rm{IM}}^{\rm{IMPM}}$ are negative, with the smallest absolute value reaching $3.86^{+0.48}_{-0.48}$. Although these absolute values are smaller than the BFs in the upper part of the Table, there is a strong evidence that the PC hypothesis is favored, because the uncertainties are low enough, implying the pipeline capability of reliably distinguishing PCs. We also note that these BFs are smaller, when comparing with the Bayesian PE of Ref. Breschi et al. (2019), performed with the time-domain full spectrum TEOBResumS-NRPM model. Nevertheless, they remains sufficiently large to robustly confirm the occurrence of a PC.
While the BF analysis demonstrates that our pipeline can detect PC signatures under favorable conditions, its reliability diminishes as the signal weakens, particularly at larger distances. Notably, for XG detectors, the SNR of the PM typically reaches values on the order of tens, when the binary is located at a distance comparable to that of GW170817 event Abbott et al. (2019a), an exceptionally favorable case. However, the SNR drops below the threshold ${\rm SNR_{thr}}\sim 7$ (see Paper II) at larger distances. As a result, detecting and characterizing the PM features becomes challenging, and even our pipeline may in principle not be able to reliably distinguish whether there is a PC to a BH or a NS remnant is formed. To address this degeneracy, we consider the alternative methods introduced in Sec. II.3 to infer PC.
We present in Fig. 13 the probability of PC computed per Eq. (12) for two equal-mass BNSs, assuming SLy EoS ($M_{\rm thr}\simeq 2.8~{}{\rm M_{\odot}}$). The first binary has a total mass $M=2.7~{}{\rm M_{\odot}}<M_{\rm thr}$ and the second has $M=3.0~{}{\rm M_{\odot}}>M_{\rm thr}$. In the former case, the probability of PC is approximately $7\%$ for IM-only and $32\%$ for IMPM. This difference comes from the different inferences of the maximum mass and the corresponding radius obtained with IM and IMPM analyses, see Fig. 10. In the IM case ${M_{\rm max}^{\rm TOV}}$ is on average larger, resulting in a higher threshold mass. In contrast, the IMPM analysis yields a more precise estimation of the maximum mass, leading to a smaller and more tightly constrained threshold mass.
For the binary with $M=3.0~{}{\rm M_{\odot}}$, the probability of PC is $80\%$, strongly favoring the PC scenario. Therefore, the first system is more likely to result in a long-lived NS remnant, while the second case supports the hypothesis of PC to a BH, as expected from the injection.
In addition, we estimate the values of the time of collapse, $t_{\rm c}$, with TEOBResumSPA_NRPMw and we compute the probability of having $t_{\rm c}>t_{\rm c}^{\rm thr}=2~{}{\rm ms}$. For the first binary ($M=2.7~{}{\rm M_{\odot}}$) we find $t_{\rm c}=23^{+15}_{-14}~{}{\rm ms}$, while for the PC injection ($M=3.0~{}{\rm M_{\odot}}$), we find $t_{\rm c}=0.28^{+0.26}_{-0.24}~{}{\rm ms}$. From Fig. 13, we notice that the probability $P\left(t_{\rm c}/t_{\rm c}^{\rm thr}>X\right)$ (green region) is vanishing for the latter binary.
Applying the same procedure to real observational data has the potential to place meaningful constraints on the properties of dense matter. Inferring the probability of PC to BH can provide valuable insights into the EoS characterization. This can help identify potential softening of the EoS, which may arise from phase transitions at extreme densities or thermal effects Radice et al. (2017); Bauswein et al. (2021); Prakash et al. (2024); Breschi et al. (2023); Fields et al. (2023).
IV Conclusion
In this paper we presented a waveform model for the complete BNS spectrum and its application to Bayesian parameter estimation with next generation detectors. We focused on single signals in Einstein Telescope’s band and the inference of neutron star matter constraints. Methodologically, our work introduces TEOBResumSPA_NRPMw, a frequency domain waveform model that describes GWs from early inspiral up to kiloHertz PM and considers injection-recovery (mock) Bayesian experiments to validate a pipeline that delivers inferences of various extreme matter constraints.
As a preparatory step to full-spectrum analyses, we first considered IM-only injections and study systematically the impact of the choice of the initial frequency (down to $2~{}{\rm Hz}$), of the signal SNR and the detector configuration on the inference. Notably, we observe an improvement of approximately one order of magnitude in the precision of the chirp mass and mass ratio when lowering the initial frequency from 20 to 2 Hz. The relative error on reduced tidal parameter improves by about a factor two. Varying the ET configuration from the triangular layout to the two-L configuration does not significantly impact the inference of the intrinsic parameters. Nevertheless, a clear improvement is seen in the measurement of the extrinsic parameters, particularly the declination, which is better constrained with the two-L configuration. This improvement is primarily due to the geographical separation of the two detectors, one in Meuse-Rhine region and the other one in Sardinia, which enhance the ability to localize the source in the sky.
Next, we focused on the full spectrum analysis with the newly developed model TEOBResumSPA_NRPMw. Our pipeline effectively recovers both the injected inspiral segment and the peak frequency of the PM signal. Other PM features are also correctly retrieved, although their constraints are less stringent compared to those on the IM and $f_{2}$. The posterior distributions of chirp mass, mass ratio and reduced tidal parameter, do not show a clear improvement when including the PM spectrum. This is attributed to the PM portion having an SNR approximately two orders of magnitude smaller than that of the IM section. Nevertheless, Bayesian model selection analysis favors the IMPM model over the IM-only, already at postmerger SNR ${\sim}8$ (where the PM detection threshold is 7). This regimes proves to be essential for probing the NS EoSs, as we demonstrate that employing the full spectrum in the analyses leads to a substantial reduction of the credibility region on the mass-radius diagram, particularly in the high-density regime.
Using an EoS sampling with minimal hypotheses (causality, validity of general relativity), we showed that a single-event PM signal plays a crucial role in constraining both the maximum density and the radius corresponding to the maximum mass, through quasi-universal relations that involves the peak frequency $f_{2}$. The inclusion of these posterior distributions significantly reduces the 90% credibility region of the mass-radius diagram, particularly at high-densities. Therefore, the possibility to perform PEs of GW signals detected with ET, that also contains PM features, enables more precise constraints on the maximum mass and the corresponding radius. In our study, we demonstrate relative errors smaller than 8% and 5% for the maximum mass and the relative neutron star radius, respectively. A novel result presented in this paper is the tight constraint on the $\Lambda(M)$ curve, where the 90% CI closely follows the theoretical model. Similarly, the inference of the pressure-density relation, especially that obtained through the IMPM analysis, yields significantly tight constraints at high-densities, again closely tracking the theoretical prediction. This approach enhances the precision in extracting information about the properties of nuclear matter under extreme conditions.
Finally, we analyze the inference of prompt collapse scenario. Bayesian model selection with IM-only and IMPM recoveries can identify PC for PM detections at or above the SNR threshold. However, other complementary approaches can be utilized. We extended the approach of Agathos et al. (2020) to infer the probability of PC. Our analysis highlights that it is possible to robustly determine whether a BNS undergoes PC or forms a NS remnant that emits PM signals, even if the latter is not detectable due to a weak PM SNR.
Our work has important limitations that should be considered for future work. First, in both IM-only and full spectrum analyses, we injected non-spinning binary systems. In the future, it would be valuable to access the ability of our pipeline to accurately recover spin components, despite the degeneracy between spin, mass ratio and tidal polarizability parameters. An ongoing study is already investigating PE without fixing any parameters in the IMPM scenario. This includes the inference of the spin components as well as additional extrinsic parameters of the binary, beyond just the luminosity distance. Moreover, our current study focuses only on the dominant (2,2) mode. Including also higher-order modes would help, for example to address the degeneracy between luminosity distance and the inclination. TEOBResumSPA already incorporates higher modes, whereas their implementation in NRPMw is still undergoing. Given that our injections are based on GW signals from XG of detectors, we must also account for the long duration of such signal. In this regime, Earth’s rotation is not negligible, and the increased sensitivity of the detector implies that multiple signals may overlap in the data stream. Consequently, future analysis techniques must address challenges in signal disentanglement and noise mitigation Abac et al. (2025); Cireddu et al. (2024); Liu et al. (2025); Wong et al. (2025). Lastly, in our analyses, we used the same model for both injection and recovery. However, Gamba et al. (2021b) showed that for signals with SNR $>80$, using different waveform models for injection and recovery introduces significant systematic errors. This highlights the importance of waveform systematics studies and motivates the development of more robust cross-model validation strategies.
Acknowledgments
The authors thank Nishkal Rao for technical contributions and discussion. GH and SB acknowledge support by the EU Horizon under ERC Consolidator Grant, no. InspiReM-101043372. SB and MB acknowledge support by the EU H2020 under ERC Starting Grant, no. BinGraSp-714626. SB acknowledges support from the Deutsche Forschungsgemeinschaft (DFG) project MEMI (BE 6301/2-1 Projektnummer: 443239082) and GROOVHY (BE 6301/5-1 Projektnummer: 523180871). MB and RG acknowledges support from the Deutsche Forschungsgemeinschaft (DFG) under Grant No. 406116891 within the Research Training Group RTG 2522/1. RG acknowledges support from NSF Grant PHY-2020275 (Network for Neutrinos, Nuclear Astrophysics, and Symmetries (N3AS)).
Computations were performed on SuperMUC-NG at the Leibniz-Rechenzentrum (LRZ) Munich and and on the national HPE Apollo Hawk at the High Performance Computing Center Stuttgart (HLRS). The authors acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at LRZ (allocations pn76li, pn36jo and pn68wi). The authors acknowledge HLRS for funding this project by providing access to the supercomputer HPE Apollo Hawk under the grant number INTRHYGUE/44215 and MAGNETIST/44288. Postprocessing and development runs were performed on the ARA cluster at Friedrich Schiller University Jena. The ARA cluster is funded in part by DFG grants INST 275/334-1 FUGG and INST 275/363-1 FUGG, and ERC Starting Grant, grant agreement no. BinGraSp-714626.
V Data availability
The waveform model developed in this work, TEOBResumSPA_NRPMw, is implemented in bajes and the software is publicly available at:
TEOBResumS is publicly available at:
Appendix A Full spectrum model
Table 3: Summary of the NR simulations used to validate TEOBResumSPA_NRPMw template. The first column reports the nature of the remnant. The following seven columns show the NR data properties, i.e.  EoS, reference, total mass $M$, mass ratio $q$, tidal polarizability $\kappa^{\rm T}_{2}$, PM peak frequency $f_{2}$ and the time of BH collapse $t_{\rm c}$.
Morph. EOS Ref. $M$ $q$ $\kappa^{\rm T}_{2}$ $f_{2}$ $t_{\rm c}$
$[{\rm M_{\odot}}]$ $[{\rm kHz}]$ $[{\rm ms}]$
Short-lived SFHo Bernuzzi et al. (2016) $2.7$ $1.0$ $79$ $3.42$ $8.0$
BLQ Prakash et al. (2021) 2.8 1.0 81 2.68 2.7
Long-lived SLy Breschi et al. (2019) $2.6$ $1.0$ $92$ $3.13$ $21$
LS220 Perego et al. (2019) 2.6 1.167 168 2.80 35
Tidally disrupted BLh Bernuzzi et al. (2020) $2.84$ $1.664$ $97$ $3.27$ $18$
DD2 Bernuzzi et al. (2024) 2.88 1.666 131 2.58 34
Figure 14: Unfaithfulness between hybrid NR-EOB waveforms and TEOBResumSPA_NRPMw model as function of the lower cut-off frequency $f_{\rm min}$. The values are computed employing ET-D Hild et al. (2011); Hild (2012) sensitivity curve and setting the upper cut-off frequency to $8~{}{\rm kHz}$. The vertical dotted lines refer to the value of the instantaneous NR frequency at merger for each case.
Figure 15: Comparison between NR data (black solid lines) and TEOBResumSPA_NRPMw model (colored dashed lines). The reported NR templates correspond to the validation set listed in Tab. 3. The time-domain TEOBResumSPA_NRPMw waveforms are estimated employing inverse FFT. For each NR simulation, we show the corresponding EoS, the mass ratio $q$ and the tidal polarizability $\kappa^{\rm T}_{2}$. The waveforms are aligned on the segment $t/M\leq-200$.
We provide here some details on the construction and the validation of the complete waveform model TEOBResumSPA_NRPMw employed in our work.
The complete model is obtained by attaching the EOB TEOBResumSPA to NRPMw at the merger. Within this framework, the parameters $\{\phi_{\rm PM},t_{0}\}$ determine the initial phase and time of the PM segment (see Paper I) and are set to the phase and time of merger from the TEOBResumSPA waveform, $\phi_{\rm mrg}$ and $t_{\rm mrg}$. In order to perform a cleaner attachment and avoid double-counting, when we employ NRPMw for the complete IMPM model, the fusion wavelet is set to zero, i.e. ${W}_{\rm fus}=0$, since the EOB model is expected to include the merger portion of data.
The validation of the complete model has been performed using 6 exemplary NR simulations from the CoRe database Gonzalez et al. (2023) computed with the thc code Bernuzzi et al. (2016); Perego et al. (2019); Bernuzzi et al. (2020); Prakash et al. (2021); Kashyap et al. (2022) and listed in Tab. 3. Note this set of simulations is not included in the calibration set. The data are chosen to cover a sufficiently large variety of phenomenologies; in particular, the validation set includes four long-lived cases, two of which with $q>1.5$ showing tidal disruption, and two short-lived cases, one of which includes deconfined quark matter. Moreover, NR data are hybridized with the time-domain TEOBResumS model, in order to extended the waveforms in the early-inspiral regime, not included in typical NR simulations. Also for these studies, we employ ET-D sensitivity curve Hild et al. (2011); Hild (2012) and we minimize over the additional set of parameters $\{{\boldsymbol{\theta}_{\rm PM}^{\rm free}},\boldsymbol{\delta}_{\rm PM}\}$ for NRPMw. Employing only the PM portion of data and NRPMw, we obtain $\log_{10}{\bar{\mathcal{F}}}$ between $-1.35$ and $-0.95$ over all the validation set of 6 binaries. In particular, short-lived cases (SFHo and BLQ) reach $\log_{10}{\bar{\mathcal{F}}}\lesssim-1.25$; while, large mass ratio binaries (BLh and DD2) recover worst matches, i.e.  $\log_{10}{\bar{\mathcal{F}}}\simeq-0.95$.
Figure 14 shows the unfaithfulness computed between TEOBResumSPA_NRPMw model and the hybrid NR-EOB waveforms, where the inner products are evaluated over the frequency range $[f_{\rm min},8~{}{\rm kHz}]$. For $f_{\rm min}<100~{}{\rm Hz}$, the model have good agreement with NR data as shown by $\log_{10}{\bar{\mathcal{F}}}\lesssim-3$, comparable to Ref. Gamba et al. (2021a). Approaching $f_{\rm mrg}$, the recovered ${\bar{\mathcal{F}}}$ shows an increasing trend, with peaks above $10^{-1}$. Subsequently, for $f_{\rm min}>1.5~{}{\rm kHz}$, the recovered ${\bar{\mathcal{F}}}$ are comparable with the values obtained only with the PM portion of data, i.e.  $\log_{10}{\bar{\mathcal{F}}}\lesssim-1$. In order to further inspect this behavior, we show in Fig. 15 the comparison between the time-domain NR data and the corresponding TEOBResumSPA_NRPMw templates, making use of the inverse FFT. The inspiral EOB template matches NR data accurately up to merger, i.e.  $t=0$; however, considerable phase differences arise after this moment in the two models. These errors can be related to the stationary-phase approximation (SPA) and to the absence of next-to-quasi-circular (NQC) corrections Riemenschneider et al. (2021); Albertini et al. (2022) in the frequency-domain EOB template. These corrections appear to be more relevant for equal-mass binaries and they motivate additional studies for the development and the calibration of accurate EOB-NR waveform models.
Appendix B Mock Bayesian inferences
In this section, we summarize the basic inference techniques implemented in bajes pipeline Breschi et al. (2021). The main novelty introduced for this work is the implementation of the relative binning technique from Dai et al. (2018); Zackay et al. (2018).
B.1 Bayesian framework
Bayesian analysis is a primary technique to perform parameter estimation. It based on the evaluation of the posterior distribution $p(\boldsymbol{\theta}|\boldsymbol{d},H)$, through Bayes’ theorem,
$p(\boldsymbol{\theta}|{\text{{d}}},H_{\rm S})=\frac{p({\text{{d}}}|\boldsymbol{\theta},H_{\rm S})p(\boldsymbol{\theta}|H_{\rm S})}{p({\text{{d}}}|H_{\rm S})},$ (16)
where the prior distribution $p(\boldsymbol{\theta}|H_{\rm S})$ represents our knowledge about the parameters we estimate, $p({\text{{d}}}|H_{\rm S})$ is the evidence and $p({\text{{d}}}|\boldsymbol{\theta},H_{\rm S})$ is called likelihood function. For a fixed set of data ${\text{{d}}}(t)$, the likelihood function is given by
$\log p({\text{{d}}}|\boldsymbol{\theta},H_{\rm S})\propto-\frac{1}{2}\sum_{i}\big{(}d-h(\boldsymbol{\theta})\big{|}d-h(\boldsymbol{\theta})\big{)}_{i}\,,$ (17)
where $H_{\rm S}$ represents the signal hypothesis, the subscript $i$ runs over the employed detectors (i.e.  it denotes that the inner products are estimated with the corresponding data series, projected waveform and PSD for each detector), and the parameter vector $\boldsymbol{\theta}$ corresponds to the combination $\boldsymbol{\theta}=\{\boldsymbol{\theta}_{\rm bin},{\boldsymbol{\theta}_{\rm PM}^{\rm free}},\boldsymbol{\delta}_{\rm fit},{\boldsymbol{\theta}_{\rm ext}}\}$ of the inspiral parameters $\boldsymbol{\theta}_{\rm bin}$, the unconstrained PM parameters ${\boldsymbol{\theta}_{\rm PM}^{\rm free}}$, the recalibration coefficients $\boldsymbol{\delta}_{\rm fit}$ and the extrinsic GW parameters ${\boldsymbol{\theta}_{\rm ext}}=\{D_{L},\iota,\alpha,\delta,\psi,t_{\rm mrg},\phi_{\rm mrg}\}$. The inner product in Eq. (17) is
[…]
with $S_{n}(f)$ the PSD of the noise segment.
Resorting to nested sampling algorithm Skilling (2006), it is possible to accurately estimate the evidence integral, i.e.
$p({\text{{d}}}|H_{\rm S})=\int p({\text{{d}}}|\boldsymbol{\theta},H_{\rm S})\,p(\boldsymbol{\theta}|H_{\rm S})\,{\rm d}\boldsymbol{\theta}\,,$ (19)
that allows us to perform model selection between two model hypotheses A and B, computing the BF of the signal hypothesis A against the signal hypothesis B as
$\mathcal{B}_{\rm{A}}^{\rm{B}}=\frac{p(d|H_{\rm B})}{p(d|H_{\rm A})}.$ (20)
When $\mathcal{B}_{\rm{A}}^{\rm{B}}>1$ ($<1$), the model hypothesis B (A) is favored. Furthermore, we introduce the SNR from Eq. (18) as
$\rho(h)=\frac{(d_{i}|h)}{\sqrt{(h|h)}}\,,$ (21)
where the definition can be extended to multiple detectors employing quadrature summation.
For a more detailed discussion on the Bayesian framework employed in GW data analysis we remand to Veitch et al. (2015); Abbott et al. (2020b); Breschi et al. (2021).
B.2 Relative binning
Relative binning Dai et al. (2018); Zackay et al. (2018); Krishna et al. (2023) is an acceleration method for PE. The most computationally expensive part of a PE is the evaluation of the likelihood function (17) with waveforms $h(f,\boldsymbol{\theta})$ in the parameter space, that are small perturbations of the best-fit waveform. Neighboring waveforms in the frequency domain are very similar to each other, therefore one can work directly with the ratio with respect to a fiducial waveform $h_{0}(f)$,
$r(f)=\dfrac{h(f)}{h_{0}(f)}.$ (22)
Since this ratio is smooth, it can be approximated by a piecewise linear function Zackay et al. (2018):
$r(f)=r_{0}(h,b)+r_{1}(h,b)(f-f_{m}(b)),$ (23)
where $b$ represents a frequency bin and $f_{m}$ is the mean frequency of the bin. Some components of the inner products involved in the likelihood computation can be precomputed to improve the efficiency:
$\begin{split}A_{0}&=\frac{4}{T}\sum_{f\in b}\dfrac{d(f)h_{0}^{*}(f)}{S_{n}},\\ A_{1}&=\frac{4}{T}\sum_{f\in b}\dfrac{d(f)h_{0}^{*}(f)}{S_{n}}(f-f_{m}(b)),\\ B_{0}&=\frac{4}{T}\sum_{f\in b}\dfrac{|h_{0}(f)|^{2}}{S_{n}},\\ B_{1}&=\frac{4}{T}\sum_{f\in b}\dfrac{|h_{0}(f)|^{2}}{S_{n}}(f-f_{m}(b))\,.\end{split}$ (24)
Thus, during the PE it is sufficient to evaluate $r_{0}(h,b)$ and $r_{1}(h,b)$ at the bin edges and compute the inner products for the likelihood as
$\begin{split}&(d|h)=\sum_{b}(A_{0}(b)r_{0}^{*}(h,b)+A_{1}(b)r_{1}^{*}(h,b)),\\ &(h|h)=\sum_{b}(B_{0}(b)|r_{0}(h,b)|^{2}+2B_{1}(b)\Re[r_{0}(h,b)r_{1}^{*}(h,b)]).\end{split}$ (25)
We then employ the binning criterion of Zackay et al. (2018) with $\epsilon=0.01$.
Relative binning technique is effective for waveform analyses up to the merger part, as the frequency domain representations are smooth in this regime. Instead, in the PM regime, the waveform structure becomes highly complex and rapidly varying in frequency, rendering relative binning inefficient and insufficiently accurate. Nevertheless, NRPMw is fast, being phenomenological, therefore acceleration techniques are not critical as in the case of TEOBResumSPA. In our full spectrum analyses, we thus employ relative binning only in the IM portion and we perform standard PE for the PM part of the spectrum. To implement this approach, we introduce cutoff frequency $f_{\rm cut}=2048~{}{\rm Hz}$. The likelihood is computed separately for the two regimes: using relative binning for frequencies below $f_{\rm cut}$ (IM) and using the full, unaccelerated Bayesian PE for frequencies above the cutoff (PM). The total likelihood is obtained by summing the two contributions, cfr. Eq (8). This hybrid approach enables efficient and accurate PE for GW signals spanning both the IM and PM regimes.
The pre-computation of some terms and the evaluation of the waveforms only on the bin edges enable PE of coalescing compact binaries in a very short amount of time: it is possible to run a PE on a single CPU in 44 minutes for a binary black holes, using the TEOBResumS template, and in 275 minutes for a BNS with TaylorF2. In the case of full spectrum BNSs waveforms with TEOBResumSPA_NRPMw, we run on a cluster with 192 CPUs in about 40 hours, making use of the MPI-parallelization of bajes Breschi et al. (2021).
Appendix C Results
We report in Table LABEL:tab:IM2 the results of the injection-recovery with IM-only TEOBResumSPA template relative to Fig. 4 and 5.
Table 4: Summary of the PEs with TEOBResumSPA for BNSs. The first eight columns report the injected parameters and the last six columns report the recovered values, with the median of the posterior distributions and the $90\%$ credibility regions.
Injected properties Recovered values
$\mathcal{M}~{}[{\rm M_{\odot}}]$ $q$ $\tilde{\Lambda}$ $\delta\tilde{\Lambda}$ $\chi_{\rm eff}$ $D_{L}~{}[{\rm Mpc}]$ $f_{0}$ SNR $\mathcal{M}~{}[{\rm M_{\odot}}]$ $q$ $\tilde{\Lambda}$ $\delta\tilde{\Lambda}$ $\chi_{\rm eff}$ $D_{L}~{}[{\rm Mpc}]$
5 2452 $1.3447868^{+0.0000005}_{-0.0000005}$ $1.212^{+0.005}_{-0.005}$ $172^{+4}_{-6}$ $40^{+30}_{-40}$ $-0.0000^{+0.0003}_{-0.0004}$ $38.6^{+0.8}_{-0.8}$
40 10 1775 $1.344787^{+0.000003}_{-0.000002}$ $1.21^{+0.02}_{-0.01}$ $171^{+4}_{-7}$ $40^{+30}_{-30}$ $0.0000^{+0.0009}_{-0.0007}$ $38.2^{+1.0}_{-0.8}$
20 1431 $1.344788^{+0.000012}_{-0.000011}$ $1.22^{+0.05}_{-0.05}$ $170^{+6}_{-12}$ $30^{+40}_{-40}$ $0.000^{+0.003}_{-0.003}$ $38.2^{+1.2}_{-1.1}$
5 1226 $1.3447868^{+0.0000010}_{-0.0000009}$ $1.213^{+0.010}_{-0.009}$ $170^{+6}_{-9}$ $40^{+30}_{-50}$ $0.0001^{+0.0005}_{-0.0004}$ $76^{+2}_{-2}$
1.3447868 1.212 172 30 0 80 10 887 $1.344788^{+0.000005}_{-0.000004}$ $1.22^{+0.03}_{-0.03}$ $168^{+7}_{-11}$ $30^{+40}_{-50}$ $0.000^{+0.002}_{-0.002}$ $75^{+3}_{-3}$
20 716 $1.34479^{+0.00002}_{-0.00002}$ $1.22^{+0.08}_{-0.08}$ $168^{+7}_{-16}$ $30^{+40}_{-50}$ $0.000^{+0.005}_{-0.004}$ $75^{+3}_{-3}$
5 817 $1.344787^{+0.0000014}_{-0.0000015}$ $1.216^{+0.014}_{-0.015}$ $167^{+8}_{-13}$ $30^{+40}_{-50}$ $0.0001^{+0.0008}_{-0.0007}$ $113^{+4}_{-4}$
120 10 592 $1.344788^{+0.000007}_{-0.000006}$ $1.22^{+0.04}_{-0.04}$ $166^{+9}_{-17}$ $30^{+40}_{-60}$ $0.000^{+0.002}_{-0.002}$ $113^{+5}_{-5}$
20 477 $1.34478^{+0.00004}_{-0.00002}$ $1.19^{+0.14}_{-0.13}$ $168^{+7}_{-18}$ $30^{+40}_{-60}$ $-0.002^{+0.008}_{-0.005}$ $112^{+6}_{-6}$
5 2391 $1.3040200^{+0.0000005}_{-0.0000005}$ $1.222^{+0.006}_{-0.005}$ $193^{+4}_{-7}$ $50^{+30}_{-40}$ $-0.0000^{+0.0003}_{-0.0003}$ $38.6^{+0.9}_{-0.9}$
40 10 1731 $1.304020^{+0.000002}_{-0.000002}$ $1.22^{+0.02}_{-0.02}$ $192^{+5}_{-8}$ $50^{+30}_{-40}$ $0.0001^{+0.0009}_{-0.0008}$ $38.2^{+1.0}_{-0.9}$
20 1396 $1.304023^{+0.000009}_{-0.000011}$ $1.24^{+0.04}_{-0.05}$ $190^{+7}_{-11}$ $40^{+40}_{-40}$ $0.001^{+0.003}_{-0.003}$ $38.2^{+1.0}_{-1.0}$
5 1195 $1.3040201^{+0.0000009}_{-0.0000009}$ $1.224^{+0.009}_{-0.009}$ $190^{+6}_{-10}$ $50^{+30}_{-50}$ $0.0001^{+0.0005}_{-0.0005}$ $76^{+2}_{-2}$
1.3040200 1.222 192 35 0 80 10 865 $1.304021^{+0.000004}_{-0.000005}$ $1.23^{+0.03}_{-0.03}$ $189^{+8}_{-14}$ $40^{+40}_{-60}$ $0.000^{+0.002}_{-0.002}$ $76^{+3}_{-3}$
20 698 $1.30402^{+0.00002}_{-0.00002}$ $1.22^{+0.09}_{-0.09}$ $189^{+7}_{-19}$ $40^{+40}_{-60}$ $-0.000^{+0.006}_{-0.004}$ $75^{+3}_{-3}$
5 797 $1.3040203^{+0.0000014}_{-0.0000013}$ $1.225^{+0.014}_{-0.013}$ $187^{+9}_{-15}$ $40^{+40}_{-60}$ $0.0002^{+0.0007}_{-0.0006}$ $114^{+5}_{-4}$
120 10 577 $1.304021^{+0.000006}_{-0.000005}$ $1.23^{+0.04}_{-0.04}$ $186^{+9}_{-14}$ $30^{+40}_{-50}$ $0.000^{+0.002}_{-0.002}$ $114^{+5}_{-6}$
20 465 $1.30402^{+0.00003}_{-0.00002}$ $1.23^{+0.13}_{-0.14}$ $186^{+10}_{-21}$ $30^{+50}_{-60}$ $0.000^{+0.008}_{-0.006}$ $113^{+6}_{-6}$
5 2344 $1.2742306^{+0.0000004}_{-0.0000004}$ $1.500^{+0.003}_{-0.003}$ $233^{+6}_{-6}$ $50^{+20}_{-20}$ $-0.0001^{+0.0003}_{-0.0005}$ $38.5^{+0.9}_{-0.7}$
40 10 1697 $1.274232^{+0.000002}_{-0.000002}$ $1.505^{+0.011}_{-0.011}$ $230^{+9}_{-10}$ $70^{+30}_{-30}$ $0.0001^{+0.0009}_{-0.0009}$ $38.1^{+1.2}_{-0.9}$
20 1369 $1.274231^{+0.000008}_{-0.000014}$ $1.50^{+0.02}_{-0.04}$ $232^{+10}_{-12}$ $80^{+20}_{-30}$ $0.000^{+0.002}_{-0.004}$ $38.1^{+1.3}_{-1.1}$
5 1172 $1.2742306^{+0.0000009}_{-0.0000010}$ $1.501^{+0.006}_{-0.007}$ $231^{+9}_{-11}$ $80^{+30}_{-40}$ $0.0001^{+0.0006}_{-0.0005}$ $76^{+2}_{-2}$
1.2742305 1.5 232 75 0 80 10 849 $1.274231^{+0.000005}_{-0.000005}$ $1.50^{+0.02}_{-0.02}$ $230^{+10}_{-20}$ $80^{+30}_{-40}$ $0.001^{+0.002}_{-0.002}$ $75^{+3}_{-3}$
20 684 $1.274230^{+0.000014}_{-0.000017}$ $1.50^{+0.04}_{-0.05}$ $230^{+10}_{-20}$ $80^{+30}_{-40}$ $0.000^{+0.004}_{-0.004}$ $75^{+3}_{-3}$
5 782 $1.2742308^{+0.0000013}_{-0.0000013}$ $1.502^{+0.009}_{-0.008}$ $230^{+10}_{-20}$ $80^{+30}_{-40}$ $0.0001^{+0.0007}_{-0.0007}$ $114^{+5}_{-5}$
120 10 566 $1.274232^{+0.000007}_{-0.000006}$ $1.51^{+0.03}_{-0.03}$ $220^{+10}_{-20}$ $70^{+30}_{-50}$ $0.000^{+0.003}_{-0.002}$ $112^{+5}_{-5}$
20 456 $1.27423^{+0.00003}_{-0.00002}$ $1.49^{+0.08}_{-0.06}$ $230^{+20}_{-30}$ $70^{+30}_{-50}$ $-0.001^{+0.007}_{-0.005}$ $111^{+6}_{-5}$
5 2238 $1.2037856^{+0.0000004}_{-0.0000004}$ $1.333^{+0.004}_{-0.004}$ $335^{+7}_{-9}$ $90^{+40}_{-40}$ $0.0000^{+0.0004}_{-0.0003}$ $38.5^{+1.1}_{-0.8}$
40 10 1620 $1.203786^{+0.000002}_{-0.000002}$ $1.334^{+0.014}_{-0.014}$ $336^{+10}_{-13}$ $90^{+50}_{-50}$ $-0.0001^{+0.0010}_{-0.0010}$ $38.3^{+1.1}_{-1.0}$
20 1307 $1.203783^{+0.000011}_{-0.000008}$ $1.32^{+0.05}_{-0.03}$ $340^{+10}_{-20}$ $90^{+50}_{-50}$ $-0.001^{+0.003}_{-0.002}$ $37.9^{+1.3}_{-1.2}$
5 1119 $1.2037856^{+0.0000008}_{-0.0000007}$ $1.333^{+0.007}_{-0.006}$ $335^{+10}_{-14}$ $100^{+40}_{-60}$ $-0.0001^{+0.0006}_{-0.0005}$ $76^{+3}_{-2}$
1.2037856 1.333 334 75 0 80 10 810 $1.203786^{+0.000004}_{-0.000005}$ $1.34^{+0.03}_{-0.03}$ $330^{+20}_{-20}$ $80^{+60}_{-80}$ $0.001^{+0.002}_{-0.00}$ $76^{+3}_{-3}$
20 654 $1.20378^{+0.00002}_{-0.00002}$ $1.33^{+0.09}_{-0.07}$ $330^{+10}_{-30}$ $90^{+50}_{-70}$ $-0.000^{+0.006}_{-0.004}$ $75^{+3}_{-3}$
5 746 $1.2037857^{+0.0000012}_{-0.0000011}$ $1.334^{+0.011}_{-0.009}$ $330^{+10}_{-20}$ $110^{+40}_{-70}$ $0.0001^{+0.0007}_{-0.0007}$ $114^{+5}_{-5}$
120 10 540 $1.203787^{+0.000006}_{-0.000006}$ $1.34^{+0.03}_{-0.04}$ $330^{+20}_{-30}$ $80^{+70}_{-80}$ $0.001^{+0.002}_{-0.002}$ $111^{+6}_{-5}$
20 436 $1.20378^{+0.00003}_{-0.00002}$ $1.33^{+0.13}_{-0.09}$ $330^{+20}_{-40}$ $80^{+60}_{-90}$ $-0.001^{+0.010}_{-0.006}$ $112^{+6}_{-6}$
5 2293 $1.2395655^{+0.0000005}_{-0.0000004}$ $1.117^{+0.009}_{-0.008}$ $385^{+4}_{-7}$ $80^{+70}_{-90}$ $0.0000^{+0.0002}_{-0.0002}$ $38.5^{+0.9}_{-0.8}$
40 10 1660 $1.239566^{+0.000002}_{-0.000002}$ $1.12^{+0.02}_{-0.02}$ $385^{+4}_{-8}$ $80^{+80}_{-90}$ $0.0001^{+0.0007}_{-0.0007}$ $38.3^{+1.0}_{-0.9}$
20 1339 $1.239564^{+0.000012}_{-0.000005}$ $1.10^{+0.08}_{-0.05}$ $385^{+4}_{-13}$ $60^{+80}_{-100}$ $-0.001^{+0.003}_{-0.001}$ $38.1^{+1.2}_{-1.0}$
5 1146 $1.2395657^{+0.0000008}_{-0.0000008}$ $1.121^{+0.015}_{-0.015}$ $381^{+7}_{-13}$ $50^{+90}_{-110}$ $0.0001^{+0.0005}_{-0.0004}$ $76^{+3}_{-2}$
1.2395655 1.117 385 35 0 80 10 830 $1.239566^{+0.000004}_{-0.000003}$ $1.12^{+0.05}_{-0.05}$ $382^{+7}_{-15}$ $50^{+90}_{-90}$ $0.000^{+0.002}_{-0.001}$ $76^{+3}_{-3}$
20 670 $1.239566^{+0.000018}_{-0.000010}$ $1.13^{+0.11}_{-0.10}$ $380^{+9}_{-24}$ $30^{+110}_{-110}$ $0.000^{+0.005}_{-0.002}$ $75^{+3}_{-3}$
5 764 $1.2395657^{+0.0000012}_{-0.0000012}$ $1.12^{+0.02}_{-0.02}$ $380^{+9}_{-19}$ $50^{+100}_{-140}$ $0.0001^{+0.0007}_{-0.0006}$ $113^{+5}_{-4}$
120 10 553 $1.239566^{+0.000006}_{-0.000005}$ $1.13^{+0.06}_{-0.08}$ $380^{+10}_{-20}$ $40^{+100}_{-120}$ $0.000^{+0.002}_{-0.002}$ $112^{+5}_{-5}$
20 446 $1.23956^{+0.00003}_{-0.00001}$ $1.12^{+0.15}_{-0.10}$ $380^{+10}_{-30}$ $30^{+120}_{-130}$ $0.000^{+0.008}_{-0.002}$ $112^{+6}_{-6}$
5 2407 $1.3153038^{+0.0000005}_{-0.0000006}$ $1.289^{+0.005}_{-0.005}$ $444^{+9}_{-9}$ $100^{+70}_{-40}$ $0.0000^{+0.0003}_{-0.0005}$ $38.5^{+0.9}_{-0.8}$
40 10 1742 $1.315303^{+0.000003}_{-0.000003}$ $1.28^{+0.02}_{-0.02}$ $449^{+10}_{-15}$ $120^{+70}_{-70}$ $-0.0004^{+0.0011}_{-0.0010}$ $38.2^{+1.1}_{-0.9}$
20 1405 $1.315296^{+0.000017}_{-0.000009}$ $1.26^{+0.07}_{-0.04}$ $452^{+9}_{-19}$ $120^{+70}_{-60}$ $-0.002^{+0.004}_{-0.002}$ $38.0^{+1.1}_{-1.0}$
5 1204 $1.3153039^{+0.0000010}_{-0.0000010}$ $1.290^{+0.009}_{-0.009}$ $444^{+13}_{-15}$ $100^{+80}_{-70}$ $-0.0001^{+0.0005}_{-0.0006}$ $76^{+3}_{-2}$
1.3153039 1.289 442 75 0 80 10 871 $1.315303^{+0.000005}_{-0.000005}$ $1.29^{+0.03}_{-0.03}$ $450^{+10}_{-20}$ $13^{+70}_{-100}$ $-0.000^{+0.002}_{-0.002}$ $75^{+3}_{-3}$
20 702 $1.315300^{+0.000024}_{-0.000019}$ $1.27^{+0.10}_{-0.08}$ $450^{+10}_{-40}$ $110^{+80}_{-100}$ $-0.001^{+0.006}_{-0.005}$ $75^{+4}_{-3}$
5 802 $1.3153038^{+0.0000016}_{-0.0000014}$ $1.29^{+0.01}_{-0.01}$ $440^{+10}_{-20}$ $120^{+70}_{-100}$ $-0.0000^{+0.0008}_{-0.0007}$ $114^{+5}_{-4}$
120 10 581 $1.315306^{+0.000006}_{-0.000008}$ $1.30^{+0.04}_{-0.05}$ $440^{+20}_{-30}$ $90^{+90}_{-110}$ $0.001^{+0.002}_{-0.003}$ $112^{+1}_{-4}$
20 468 $1.31530^{+0.00003}_{-0.00002}$ $1.27^{+0.11}_{-0.10}$ $445^{+15}_{-30}$ $110^{+80}_{-110}$ $-0.001^{+0.008}_{-0.006}$ $112^{+6}_{-5}$
5 2268 $1.2236917^{+0.0000005}_{-0.0000005}$ $1.286^{+0.005}_{-0.005}$ $530^{+9}_{-11}$ $150^{+70}_{-60}$ $-0.0000^{+0.0004}_{-0.0006}$ $38.6^{+1.0}_{-0.9}$
40 10 1642 $1.223691^{+0.000003}_{-0.000003}$ $1.28^{+0.02}_{-0.02}$ $535^{+7}_{-15}$ $190^{+40}_{-90}$ $-0.0005^{+0.0012}_{-0.0011}$ $38.2^{+1.2}_{-0.9}$
20 1325 $1.223690^{+0.000008}_{-0.000008}$ $1.28^{+0.04}_{-0.04}$ $533^{+9}_{-17}$ $160^{+60}_{-80}$ $-0.001^{+0.003}_{-0.002}$ $38.1^{+1.3}_{-1.2}$
5 1134 $1.2236918^{+0.0000009}_{-0.0000008}$ $1.287^{+0.009}_{-0.008}$ $529^{+11}_{-17}$ $160^{+60}_{-100}$ $-0.0000^{+0.0006}_{-0.0005}$ $76^{+2}_{-2}$
1.2236919 1.288 526 110 0 80 10 821 $1.223692^{+0.000004}_{-0.000004}$ $1.29^{+0.03}_{-0.03}$ $529^{+13}_{-21}$ $170^{+60}_{-100}$ $-0.000^{+0.002}_{-0.002}$ $75^{+3}_{-3}$
20 662 $1.22369^{+0.00002}_{-0.00001}$ $1.30^{+0.09}_{-0.07}$ $522^{+18}_{-38}$ $130^{+90}_{-100}$ $0.001^{+0.006}_{-0.004}$ $75^{+3}_{-3}$
5 756 $1.2236919^{+0.0000014}_{-0.0000012}$ $1.289^{+0.014}_{-0.012}$ $525^{+14}_{-26}$ $160^{+70}_{-130}$ $0.0001^{+0.0007}_{-0.0006}$ $113^{+5}_{-4}$
120 10 547 $1.223694^{+0.000006}_{-0.000007}$ $1.30^{+0.04}_{-0.05}$ $520^{+20}_{-33}$ $140^{+90}_{-120}$ $0.001^{+0.003}_{-0.003}$ $112^{+6}_{-5}$
20 442 $1.22369^{+0.00003}_{-0.00002}$ $1.29^{+0.12}_{-0.11}$ $523^{+18}_{-48}$ $140^{+80}_{-140}$ $0.000^{+0.009}_{-0.007}$ $112^{+5}_{-6}$
5 2089 $1.1076460^{+0.0000004}_{-0.0000003}$ $1.125^{+0.008}_{-0.009}$ $570^{+6}_{-10}$ $90^{+140}_{-110}$ $0.0000^{+0.0003}_{-0.0003}$ $38.4^{+0.8}_{-0.8}$
40 10 1514 $1.107646^{+0.000002}_{-0.000002}$ $1.13^{+0.03}_{-0.03}$ $569^{+6}_{-16}$ $60^{+100}_{-100}$ $0.0002^{+0.0008}_{-0.0008}$ $38.2^{+1.4}_{-1.0}$
20 1222 $1.107646^{+0.000008}_{-0.000006}$ $1.13^{+0.07}_{-0.06}$ $567^{+7}_{-17}$ $50^{+120}_{-90}$ $0.000^{+0.003}_{-0.002}$ $38.0^{+1.3}_{-1.1}$
5 1045 $1.1076460^{+0.0000007}_{-0.0000007}$ $1.126^{+0.014}_{-0.014}$ $567^{+8}_{-12}$ $140^{+90}_{-160}$ $0.0001^{+0.0005}_{-0.0005}$ $76^{+2}_{-3}$
1.1076460 1.125 570 50 0 80 10 757 $1.107647^{+0.000003}_{-0.000002}$ $1.13^{+0.04}_{-0.03}$ $568^{+9}_{-29}$ $80^{+160}_{-260}$ $0.0003^{+0.0016}_{-0.0010}$ $76^{+3}_{-3}$
20 611 $1.107646^{+0.000017}_{-0.000008}$ $1.14^{+0.12}_{-0.10}$ $565^{+10}_{-32}$ $70^{+150}_{-150}$ $0.000^{+0.006}_{-0.003}$ $75^{+3}_{-3}$
5 696 $1.1076461^{+0.0000011}_{-0.0000009}$ $1.13^{+0.02}_{-0.02}$ $564^{+11}_{-28}$ $90^{+130}_{-220}$ $0.0001^{+0.0006}_{-0.0006}$ $113^{+5}_{-5}$
120 10 505 $1.107646^{+0.000005}_{-0.000005}$ $1.13^{+0.06}_{-0.09}$ $565^{+11}_{-25}$ $90^{+140}_{-220}$ $0.000^{+0.002}_{-0.002}$ $113^{+5}_{-5}$
20 407 $1.107643^{+0.000019}_{-0.000007}$ $1.11^{+0.14}_{-0.09}$ $560^{+14}_{-39}$ $90^{+190}_{-170}$ $-0.001^{+0.007}_{-0.002}$ $112^{+6}_{-6}$